### v1

# HAMLET: Hyperadaptive Agent-based Modeling for Live Embodied Theatrics

Abstract
Creating an immersive and interactive theatrical experience is a long-term goal in the field of interactive narrative. The emergence of large language model (LLM) is providing a new path to achieve this goal. However, existing LLM-based drama generation methods often result in AI agents that lack initiative and cannot interact with the physical environment. Furthermore, these methods typically require detailed user input to drive the drama. These limitations reduce the interactivity and immersion of online real-time performance. To address the above challenges, we propose HAMLET, a multi-agent framework focused on drama creation and online performance. Given a simple topic, the framework generates a narrative blueprint, guiding the subsequent improvisational performance. During the online performance, each actor is given an autonomous mind. This means that actors can make independent decisions based on their own background, goals, and emotional state. In addition to conversations with other actors, their decisions can also change the state of scene props through actions such as opening a letter or picking up a weapon. The change is then broadcast to other related actors, updating what they know and care about, which in turn influences their next action. To evaluate the quality of drama performance, we designed an evaluation method to assess three primary aspects, including character performance, narrative quality, and interaction experience. The experimental evaluation shows that HAMLET can create expressive and coherent theatrical experiences. Our code, dataset and models are available at https://github.com/HAMLET-2025/HAMLET.
Figure 1: The HAMLET framework creates AI drama in two main stages. First, during offline planning, a collaborative workflow of agents including the actor designer, plot designer, and reviewer creates initial materials, which are then integrated by a director agent into a structured narrative blueprint. This blueprint then guides the subsequent online performance, where a control system composed of a planner, transfer, and advancer directs a dynamic and improvisational theatrical experience. The figure illustrates this entire process using the play-within-a-play scene from Hamlet, Act III, Scene II.
Introduction
In recent years, large language models (LLMs) have demonstrated strong ability in various generative tasks, especially in creative fields such as story creation (Li et al. 2024) and role-playing (Tu et al. 2024). These models are able to generate fluent and imaginative texts, providing a foundation for the development of interactive narrative. When applying these capabilities to the creation of drama, the exploration can be divided into two parts: drama generation and drama performance.
Existing studies have explored drama generation using language models, employing methods such as hierarchical approaches (Fan, Lewis, and Dauphin 2018; Yao et al. 2019) and the modular generation of script structures (Mirowski et al. 2023). However, significant challenges exist in applying LLMs to structured drama generation, especially in scenarios requiring online real-time enactment. In interactive drama performance, LLM agents take on the role of performing the narrative. In contrast to free-form story creation, drama requires each actor to make decisions and take actions consistent with their profiles and the overall plot progression. These interactions must unfold across a series of scenes to collaboratively advance the narrative. Furthermore, existing methods typically require detailed user input, such as a full story outline (Wu et al. 2024) or elaborate guiding paragraphs (Wu et al. 2025c), which not only increases the design cost, but also reduces the possibility of an arbitrary storyline.
The challenge of the drama performance is to redefine the interaction patterns of LLM agents. In many prior studies, these patterns are limited to passive, turn-based responses. For example, some studies focus on the adequacy of single-turn dialogues rather than active participation (Wu et al. 2024). AI agents typically await user instructions, lacking genuine proactivity. This reactive model is hard to support dynamic scenes where multiple AI actors interact with human players. We believe that truly vibrant, dramatic interactions require AI actors to make autonomous decisions, collaborate or conflict in open scenarios, and actively guide the development of the plot. This paradigm shift from passive response to active guidance is the core advocated by Agentic AI (Sapkota, Roumeliotis, and Karkee 2025), providing a new perspective on how to enhance actor initiative in dramatic performance.
In addition to the interaction of agents, a comprehensive framework for drama performance should also incorporate two other critical aspects: physical environment interaction and an overall evaluation system. First, drama is an art of both language and action. Actors’ behaviors should affect their physical environment, and the environment’s feedback is a vital part of the performance. Lacking a physical setting, the drama feels more like an abstract dialogue than a tangible, immersive world. Second, there is currently no effective evaluation method to measure the quality of the drama online performance. Most existing LLM benchmarks focus on text generation quality (Gómez-Rodríguez and Williams 2023) or role-playing ability (Tu et al. 2024; Wu et al. 2025a; Wang et al. 2025), rather than the comprehensive effect of the entire drama performance.
To address the above challenges, we propose HAMLET, a multi-agent framework that enables Hyperadaptive Agent-based Modeling for Live Embodied Theatrics. Specifically, we make the following contributions:
  • A HAMLET framework: Our framework is designed as a multi-agent collaborative workflow, divided into two key stages: offline planning and online performance. Offline planning only needs a simple topic to generate a structured narrative blueprint. This blueprint provides freedom for improvisation of online performance while ensuring the integrity of the main story structure. To enhance the online performance, a Perceive And Decide (PAD) module acts as a pre-response module for each agent, making their perceptions and decisions more human-like.
  • A comprehensive evaluation method and leaderboard: To objectively assess the quality of drama generation and performance, we establish a comprehensive evaluation method that assesses three key dimensions – character performance, narrative quality, and interaction experience. This method incorporates a leaderboard that uses GPT-4o as a strong baseline for win rate comparisons. Additionally, we trained HAMLETJudge, a critic model designed for cost-effective and reliable drama performance evaluation.
  • Extensive experiments: We conducted evaluations of HAMLET’s core components. Notably, both PAD and HAMLETJudge, are 8B models, yet they achieved state-of-the-art performance in their respective tasks.
Related Work
LLM-Based Drama
Research on LLM-based drama has evolved along two main directions: drama generation and drama performance. The initial efforts in drama generation adapted techniques from general story creation, using hierarchical models to plan plots and generate coherent narratives (Fan, Lewis, and Dauphin 2018; Yao et al. 2019). With the development of this field, researchers have tried multi-LLM collaboration (Venkatraman, Tripto, and Lee 2024) and hierarchical methods (Mirowski et al. 2023) that separate planning from generation. A key limitation of these systems is that they require a complete story (Wu et al. 2024) or a leading paragraph (Wu et al. 2025c) as input. Unlike previous work, our work presents a system designed for arbitrary, on-demand drama generation. At the same time, the research focus has also expanded to the second dimension: drama performance. Role-playing in drama performance is usually achieved by LLM agents (Chen et al. 2024b). These agents require knowledge about the actor to learn its expression style (Shanahan, McDonell, and Reynolds 2023; Wang et al. 2023; Tu et al. 2024). Current role-playing systems are built using either prompt engineering (Li et al. 2023; Chen et al. 2023) or fine-tuning (Shao et al. 2023; Lu et al. 2024). However, these methods find it hard to capture complex character traits from simple descriptions (Huang et al. 2023). To create more natural actors, better models are needed. We propose a perceive and decide model to improve dialogue consistency and emotional expression.
Evaluation for Role-Playing Conversation Agents
Traditional dialogue metrics like Embedding Similarity, BLEU, and ROUGE (Mou et al. 2016; Wu et al. 2020; Serban et al. 2017) are widely used in role-playing conversation agent (RPCA) research (Wang et al. 2024; Zhou et al. 2024; Tao et al. 2024). However, these quantitative indicators struggle to evaluate role consistency and personality. To address this, specialised evaluation frameworks have been proposed. RoleEval (Shen et al. 2023) uses role-specific multiple-choice questions to test the model’s understanding of a character. SocialBench (Chen et al. 2024a) constructs evaluation questions from multi-source dialogues, while CharacterEval (Tu et al. 2024) adopts multi-round dialogues and multi-dimensional scoring for aspects like conversational ability. Additionally, RAIDEN (Wu et al. 2025b) builds a Q&A dataset via annotator interaction to evaluate responsiveness in specific dimensions. Although these methods provide quantitative criteria, they are typically limited to two-character or user-character scenarios. While CoSER (Wang et al. 2025) expands the number of roles, it still lacks an evaluation mechanism for overall dramatic performance. Building on this previous work, our study proposes new dimensions to evaluate the complete dramatic performance.
Figure 2: An illustration of HAMLET’s core components for performance generation: a narrative blueprint that defines the scene, plot and character profiles, and the resulting real-time conversation containing scene descriptions and dialogue.
Figure 3: An example of the real-time interaction and adjudication loop in the online performance. An actor agent attempts an action or speech, termed a beat, to progress towards the current narrative point. The narrator agent then intercepts this attempt, determines whether it is a success or failure, and provides objective feedback to all participants in the drama.
HAMLET: A Multi-Agent Framework Design for AI Drama
To achieve an automated, interactive, and expressive AI drama experience, we designed the HAMLET framework, as shown in Figure 1. The framework decouples the generation and performance of the drama into two main stages: offline planning, a multi-agent collaborative workflow to generate a narrative blueprint, and online performance, a hierarchical control system that executes the blueprint, manages real-time interactions, and handles environmental feedback.
Offline Planning
The target of the offline planning stage is to transform user input into a structured narrative blueprint. This stage is designed to handle two types of input: 1) An arbitrary and customizable topic. For this input type, the offline workflow directly generates the content for a complete act based on the provided topic. 2) A complete literary work. When given a full literary text, the workflow first deconstructs it into a series of acts based on its chapters and content structure. Subsequently, it proceeds with drama design for each act, ensuring the adaptation remains faithful to the source material.
Agent-Based Workflow Architecture
Our offline workflow consists of four agents: the actor designer, the plot designer, the reviewer, and the director. This workflow implements character creation, plot generation, review, and final integration.
Character Profile Generation and Review
The starting point of the workflow is actor construction. The actor designer is responsible for generating actor profiles of core characters based on user input. To ensure the depth and accuracy of character creation, this agent queries an external knowledge base via its search module. The output is a structured actor profile that defines the character’s static attributes, such as background and personality, and dynamic attributes, such as their initial goal and core relationships within the story. The profile is then submitted to the reviewer, who checks the rationality of the character settings, the clarity of motivations, and the relationships between actors.
Plot Generation and Structuring
After all actor profiles are approved, the plot designer composes a preliminary narrative draft based on the topic and actors. This draft is then submitted to the reviewer for evaluation. Once the review is passed, the director is responsible for the final structural processing, reconstructing the linear story draft into a hierarchical plot profile. This process includes the following key steps: 1) Act and scene definition: Divide the drama into several acts and specify the scenes in which each act takes place. 2) Creation of environmental elements: Generate a list of interactive props for each scene. This list has specific descriptions and location information for the props. 3) Definition of points: In each act, define a series of narrative points. Each point contains a clear flag and a result that marks its completion. 4) Backward planning: In order to prevent the plot from deviating, the director will give priority to generating the end point when planning an act. Then, based on this end point, it will supplement and construct the logically coherent preceding points leading to the ending in a reverse manner.
Finally, the director integrates the plot profile with the actor profile to generate a narrative blueprint which is shown in Figure 2(a) as the input of the online performance stage.
Online Performance
Upon receiving the narrative blueprint, the online performance stage begins to transform it from a static plan into a dynamic, interactive, and immersive environment. This environment will accommodate both autonomous AI actors and human players who assume certain roles. To ensure the performance effect, this phase introduces more specific narrative units, environmental interaction mechanisms, and a group of collaborative agents. An example of online real-time performance is illustrated in Figure 2(b).
Performing Drama
Online performances are based on acts. Each act consists of two components: scene and point. An act can contain one or more scenes and points, which together cover the complete drama process of the act. The scenes define a physical environment in which the drama takes place. It is the stage of ”where” and contains all interactive props. The points define the goals of the plot and are milestones of ”what to do”.
The narrative path between two points is formed by a dynamically generated trajectory which consists of a series of beats. We define a beat as an effective interaction step in which an actor takes an effective action. This action logic is driven by a dual-goal system. For each beat, an actor’s decision references both the public flag of the current point and a personal private goal that is refreshed at the beginning of every act.
Due to the actors’ autonomy, multiple trajectories can connect $point_{i}$ to $point_{i+1}$, such as one composed of $beat_{j_{1}},beat_{j_{2}},\dots$ and another of $beat_{k_{1}},beat_{k_{2}},\dots$. These multiple combinations introduce a high degree of freedom and arbitrariness to the performance.
Figure 4: The perceive and decide module processes external stimulus and internal state to determine a response strategy by tool calling.
Environment Interaction
Credible physical interaction is the core challenge of the online performance stage. As shown in the Figure 3, we designed a narrator agent to adjudicate all interactions between actors and the environment. The role of the narrator is to ensure the rationality of all physical actions. When an actor attempts to perform a physical action, the narrator will make a judgment based on the environment state and physical rules. If the action is feasible, the narrator will confirm the success, update the environment state, and broadcast an objective description to all participants. Otherwise, the narrator will determine failure with a logical explanation.
Perceive and Decide Module
Hierarchical Control of Performance Effects
All AI actors in the drama performance utilize a layered architecture, which is comprised of an LLM and a PAD module. The LLM is responsible for generating specific speech and actions, while the PAD handles the strategic decision-making that guides them. The design of the PAD is detailed in the following subsection.
To ensure the online performance balances improvisational freedom with stable progression along the main storyline, we designed three collaborative control agents: 1) Planner pre-designs and reviews multi-trajectory result. It breaks down the flag into a few sequence of executable beats which forms a trajectory. Beats can be references for advancer to guide the plot forward. 2) Transfer moves the story to the next point by regularly checking if the flag is met. Once satisfied, it advances the drama to the next point and manages relevant actors to enter or leave. 3) Advancer ensures the plot progresses. If it stalls beyond a set time threshold, the advancer directs necessary actors based on the current flag or next beat. Their relationship is illustrated in Figure 3.
In the HAMLET framework, the PAD module plays a crucial role in guiding the AI actors to generate final responses. The design of PAD is conceptually grounded in the dual-process theory of human cognition proposed by Kahneman (Kahneman 2011). According to this theory, human thinking is characterized by two distinct modes of processing: System 1, which is fast, automatic, and intuitive; and System 2, which is slower, deliberate, and analytical. PAD integrates both intuitive and reflective reasoning mechanisms to better simulate human-like decision-making in complex and nuanced drama contexts. Specifically, PAD is mainly responsible for generating a decision of fast, slow, silence or potential actions by tool calls—to simulate and extend the dual-system mechanism. This expanded set of response modes enhances the module’s compatibility with the interactive real-time dramatic context proposed in the study.
Design Principles
As depicted in Figure 4, the core input of PAD is based on a dual principle: subjective and objective perspectives. When perceiving, PAD acts as an AI actor possessing both subjective awareness of its internal state and an objective perception of external stimulus. To ensure well-grounded decision-making, PAD integrates information from both perspectives when a new event occurs.
The first perspective is the internal state, representing the actor’s self-awareness. This is maintained as a live profile containing static components Persona and Subjective Relationships as well as dynamic retrieved or updated elements like Memory and Goal.
The second perspective addresses external stimulus, which constitutes the objective environmental and contextual information. The external stimulus is composed of Environment Description, Actor List, Dialogue History and Interactable Objects. Together, these components form a comprehensive representation of the external context, facilitating the integration of situational awareness into the decision-making process.
Algorithm 1 Perceive and Decide Model Working Process
1:Actor $a_{k}$, Dramatic Context $\mathcal{C}_{drama}$, PAD Model Parameters $\pi_{\theta}$
2:A final decision tuple $(r,o)$, where $o$ is the structured action (composed of subject, verb, and object) or empty ($\varnothing$), $r\in\{\texttt{FAST},\texttt{SLOW},\texttt{SILENCE}\}$ is the response strategy.
3:procedure GetActorDecision($a_{k},\mathcal{C}_{drama},\pi_{\theta}$)
4:     $prompt_{pad}\leftarrow\mathrm{EncodeStrategy}(a_{k},\mathcal{C}_{drama})$
5:     $\hat{r},[\hat{thinking}]\leftarrow\mathrm{PAD}(prompt_{pad},\pi_{\theta})$
6:     $r\leftarrow\arg\max_{r^{\prime}\in\{\dots\}}P(r^{\prime}|\hat{r})$
7:     $\hat{o}\leftarrow\mathrm{GenerateAction}(r,[\hat{thinking}],a_{k},\mathcal{C}_{drama})$
8:     $o\leftarrow\mathrm{ParseAction}(\hat{o})$
9:     return $(r,o)$
10:end procedure
Decision-making Process
As detailed in algorithm 1, the decision-making process translates abstract strategic intent into concrete, executable actions through a two-stage process. First, PAD determines a high-level response strategy to set the reaction’s timing and tone. It can also generate the underlying thinking for this choice as an internal monologue through a chain-of-thought process. Subsequently, both the strategy and the generated thinking are used to guide an LLM, which acts as a tool to produce the final, concrete behavior. This behavior is composed of the specific speech to be delivered and a structured action, which is parsed into an executable triplet.
Evaluation Method
Dataset Construction
Our public evaluation dataset was constructed from 100 diverse cases, comprising 50 literary excerpts and 50 custom-authored drama topics. The literary excerpts are selections from 25 classical Chinese [^1] https://m.douban.com/subject_collection/ECKM5FBEI and 25 renowned English [^2] https://sites.prh.com/modern-library-top-100/#top-100-novels literary works. The 50 custom topics, designed to span 10 distinct themes, were developed and reviewed carefully by our annotators. Figure 5 illustrates the thematic distribution of the entire dataset. More details about the public dataset can be found in Appendix A.
Figure 5: Distribution of drama topics across the dataset.
Model English Chinese Overall Score
Character Narrative Interaction Average Character Narrative Interaction Average
Non-Reasoning Models
Claude-4-sonnet 76.50 77.30 76.96 76.92 80.18 79.20 79.66 79.68 78.30
Claude-3.7-sonnet 65.80 66.94 65.98 66.24 76.00 75.20 75.48 75.56 70.90
Claude-3.5-sonnet 61.00 62.15 62.10 61.75 60.50 59.00 59.99 59.83 60.79
Gemini-2.5-flash 46.00 47.10 46.70 46.60 52.00 52.90 52.15 52.35 49.48
GPT-4.5-preview 59.21 60.00 59.92 59.71 61.50 62.50 61.91 61.97 60.84
GPT-4.1 63.50 62.49 62.98 62.99 62.00 62.80 62.64 62.48 62.74
Mistral-Small-3.2-24B 48.00 48.50 48.22 48.24 51.00 51.83 52.00 51.61 49.93
DeepSeek-V3-0324 54.00 55.13 55.00 54.71 64.50 64.00 64.01 64.17 59.44
Llama-3.1-70B 49.80 49.00 48.86 49.22 42.00 42.81 42.00 42.27 45.75
Llama-3.1-8B 35.00 36.01 35.52 35.51 33.50 34.00 33.99 33.83 34.67
Higgs-Llama-3-70B 72.00 78.50 66.22 72.24 64.00 64.50 63.95 64.15 68.20
Qwen3-8B 47.50 46.80 47.00 47.10 58.00 57.50 58.05 57.85 52.48
Qwen3-32B 65.00 65.88 65.20 65.36 66.00 65.50 65.84 65.78 65.57
Qwen3-235B-A22B 69.50 69.80 69.65 69.65 72.50 73.00 72.78 72.76 71.21
Reasoning Models
Gemini-2.5-pro 61.00 62.22 61.70 61.64 62.00 62.80 62.25 62.35 62.00
Claude-4-sonnet-Thinking 79.50 78.40 79.04 78.98 78.42 80.32 81.03 79.92 79.45
Minimax-M1 51.50 52.32 52.00 51.94 65.00 65.50 65.19 65.23 58.59
OpenAI-o3 69.00 69.95 69.40 69.45 78.00 77.50 78.17 77.89 73.67
Magistral-Small-2506 59.00 60.00 59.74 59.58 58.50 59.30 58.90 58.90 59.24
DeepSeek-R1-0528 66.00 67.10 66.64 66.58 79.00 79.50 79.61 79.37 72.98
Qwen3-8B-Thinking 50.00 51.61 51.00 50.87 65.80 65.00 65.55 65.45 58.16
Qwen3-32B-Thinking 69.50 68.80 69.00 69.10 78.00 79.00 78.77 78.59 73.85
Qwen3-235B-A22B-Thinking 70.50 71.00 70.72 70.74 76.00 75.80 75.96 75.92 73.33
Table 1: Performance evaluation of different models based on HAMLETJudge. Bold values indicate the best performance in each column.
Dimension Definition
Defining evaluation dimensions is crucial for quantifying the performance of real-time drama generation. We categorize our evaluation into three core dimensions: Character Performance, Narrative Quality, and Interaction Experience.
Character Performance (CP) measures the quality of the AI actors. This dimension evaluates Believability, i.e., the consistency of characters with their established personas, and Agency, which reflects the richness of emotional expression and the character’s ability to effectively advance the narrative.
Narrative Quality (NQ) examines the story’s overall craftsmanship. This includes its Coherence, ensuring the logical development of the plot; its Resonance, measured by its thematic relevance and depth; and its Integrity, which evaluates the structural completeness of the storyline from beginning to end.
Interaction Experience (IE) focuses on the AI actor’s engagement with the system. This dimension encompasses the quality and timeliness of the system’s reactions, termed Responsiveness; the level of cognitive and emotional engagement, or Immersion; and the overall technical smoothness of the interaction, referred to as Fluency.
Evaluation Principle
Our evaluation principle is to holistically assess dramatic performance, rather than evaluating agent responses individually, turn by turn. This approach is crucial because drama often incorporates literary devices such as foreshadowing and plot twists. Consequently, a seemingly subpar or unusual generation in a single turn might be a deliberate narrative setup for subsequent developments and should not be penalized in isolation.
For the whole drama performance result, we employ HAMLETJudge to automate the evaluation process. This model conducts pairwise comparisons between test result and the baseline result, assigning a score based on a 5-point Likert scale (Robinson 2014) to determine a win rate. The scoring guideline is defined in Appendix B.
Experiments
In this section, we present the experimental results to demonstrate the superiority of our proposed HAMLET framework, alongside ablation studies verifying its reliability and validity.
Settings
To ensure rigorous and reproducible evaluation, we define the baseline and test configurations clearly. All underlying models except the PAD component in HAMLET share the same GPT-4o backbone, with a greedy sampling strategy. Regarding user settings, each agent can be freely configured, and the PAD component is optional.
HAMLET Leaderboard
Following the settings above, we compared various mainstream LLMs ranging from open-source to closed-source and non-reasoning to reasoning models. This reveals their capabilities in English and Chinese online drama performance, serving as a practical reference for real-world applications. The results are shown in Table 1.
Reliability and Validity of HAMLET
Reliability
HAMLETJudge was fine-tuned on pairwise data labeled by annotators. To create this dataset, annotators manually compared paired results and provided judgments with explanations. To validate the effectiveness of our judge model, we measured its agreement with a held-out human validation set using the Pearson correlation coefficient. As shown in Table 2, HAMLETJudge closely aligns with human assessments, outperforming other strong models such as GPT4.1, Claude-4-sonnet and Gemini-2.5-pro.
Metric HamletJudge GPT-4.1 Claude-4-Sonnet Gemini-2.5-Pro
CP 0.792 0.675 0.698 0.720
NQ 0.807 0.593 0.783 0.684
IE 0.773 0.622 0.804 0.701
Average 0.791 0.630 0.762 0.702
Table 2: Pearson correlation coefficients (Pearson 1901) of different models and human evaluators.
To further demonstrate the reliability of PAD, we conducted additional experiments evaluating model performance under various response strategies. It is worth noting that, models with stronger general ability tend to have a more homogeneous and consistent performance. Models with output reasoning tokens before tool calls may cause an unbearable time delay in real-time drama. A latency penalty is introduced to capture this effect. The experiment result is shown in Table 3.
Our analysis reveals a distinct performance-latency trade-off among existing models. Models engineered for explicit reasoning achieve balanced performance across strategies but suffer from significant latency penalties. Conversely, non-reasoning models are faster but exhibit a strong performance bias, excelling in Fast scenarios while lacking robustness in complex interactions. In contrast, PAD resolves this problem. It achieves the highest final score by delivering top performance across all strategies, all with zero latency.
Model (FC) Responding Strategy Latency Penalty Final Score
Fast Slow Silence
Qwen2.5-7B-Instruct 0.779 0.359 0.131 0 0.423
Qwen2.5-72B-Instruct 0.692 0.452 0.262 0 0.474
Qwen3-8B 0.699 0.452 0.066 0 0.406
Qwen3-32B 0.773 0.396 0.098 0 0.422
Qwen3-8B-Thinking 0.668 0.321 0.459 0.05 0.532
Qwen3-32B-Thinking 0.668 0.434 0.361 0.10 0.538
GPT-4o 0.556 0.623 0.328 0 0.502
GPT-4.1-mini 0.909 0.132 0.180 0 0.407
DeepSeek-R1-0528 0.723 0.615 0.470 0.15 0.453
Gemini-2.5-pro 0.742 0.536 0.519 0.10 0.449
PAD 0.822 0.736 0.711 0 0.756
Table 3: Evaluation under responding strategies by tool calling. Underlined values reflect an applied latency penalty.
Validity
In this ablation study, we assess the impact of core design choices by comparing three distinct experimental setups. We randomly select 30 topics from the public evaluation dataset and control the experiment settings as GPT-4o under greedy strategy, then compare: drama performed by (i) raw-prompted GPT-4o only; (ii) the multi-agent collaboration of the full HAMLET framework; and (iii) HAMLET framework with PAD disabled.
Figure 6: Ablation study of HAMLET framework design.
As shown in Figure 6, prompting a single GPT-4o yields substantially lower performance, highlighting the necessity of our multi-agent workflow design. Furthermore, PAD-enabled HAMLET consistently outperforms the version with PAD disabled in all 10 topic categories, demonstrating that PAD serves as a crucial component in the online drama rendition, making the interaction and conversation of AI actors more natural, coherent and humanistic.
Case Study
To further demonstrate the working mechanism of the HAMLET framework and each of its component designs, we select some real cases as shown in Table 6 in Appendix C. In Case 1, the narrator first considers the situation before responding to make a reasonable judgment about the AI actor’s input. Our framework also supports users taking on a certain character role (Cases 2, 3, and 4), where the narrator, transfer, and advancer can handle various circumstances (such as non-existent props, irresponsible actions, and stubborn choices) to ensure the drama runs smoothly. On the other hand, the planner is responsible for multi-trajectory design and review: the same goal can be met by different beat trajectories as long as the process is reasonable (Cases 5 and 6).
Annotation Statement
This work relies on human annotation, which is critical for both the training and evaluation of our proposed models. The annotation of PAD and HAMLETJudge datasets was conducted by a team of five human experts, each with at least one year of professional experience in relevant fields such as creative writing or drama. To ensure high-quality annotations, we provided clear and task-specific instructions to all labelers. In addition, the datasets underwent rigorous safety checks to ensure no personally identifiable information or potentially problematic content is available. The datasets will be released under the CC BY-NC 4.0 license. The labeling instructions can be found in Appendix E and F.
Conclusion
In this paper, we introduced HAMLET, a multi-agent framework designed to address key challenges in AI-driven theatre. HAMLET generates a guiding narrative blueprint from a simple topic and provides agents with a perceive and decide module, enabling autonomous thinking and physical environmental interaction. To assess performance, we established a comprehensive evaluation method and a leaderboard, where our specialized critic model, HAMLETJudge, achieved top-ranking results. Extensive experiments show that our approach creates expressive and coherent theatrical experiences, paving a new path toward autonomous and immersive interactive drama.
Appendix A A. Detailed information of 100 Cases Selected for Public Dataset
This appendix details the public dataset of 100 cases in Table 4. The dataset comprises three parts: 1) excerpts from 25 Chinese literary works, 2) excerpts from 25 English literary works, and 3) 50 custom-authored drama topics. For the HAMLET Leaderboard evaluation result in Table 1, the score for both English and Chinese is calculated on a 75-item set, which consists of the 25 literary excerpts and 50 custom topics in the corresponding language.
Selected established literature workpieces
1. Dream of the Red Chamber 2. Journey to the West
3. Romance of the Three Kingdoms 4. Water Margin
5. The Three-Body Problem 6. To Live
7. Four Generations Under One Roof 8. Memories of Peking: South Side Stories
9. Demi-Gods and Semi-Devils 10. The Smiling, Proud Wanderer
11. Wandering 12. Rickshaw Boy
13. Straw House 14. The Bronze Age
15. Border Town 16. The Chess Master
17. The Golden Age 18. My Father’s Back
19. Blossoms 20. Frog
21. Cat Country 22. That Unknown Story
23. Farewell My Concubine 24. White Deer Plain
25. Fortress Besieged (Items 1–25 are translated titles of Chinese literary works.) 26. One Hundred Years of Solitude (Items 26–50 are original titles of English literary works.)
27. Brave New World 28. A Clockwork Orange
29. The Time Traveler’s Wife 30. The Princess Bride
31. The Secret Garden 32. The Outsiders
33. The Call of the wild 34. Little Women
35. Hamlet 36. The Odyssey
37. Harry Potter 38. Frankenstein
39. The Kite Runner 40. King Lear
41. The tragedy of Macbeth 42. The Adventures of Huckleberry Finn
43. Life of Pi 44. A Tale of Two Cities
45. The tempest 46. Romeo and Juliet
47. The Adventures of Sherlock Holmes 48. Wuthering Heights
49. Catch-22 50. Don Quixote
Costomizable drama topic design
51. Porco Rosso and Gina discuss topics about war, love and responsibility in a café, and after a while Phil also arrives.
52. Kenshin Himura, the wandering swordsman, walked into the café carrying his reverse-blade sword, only to find his late wife, Tomoe Yukishiro—who had died years ago saving him—standing there.
53. Conan and Gin engaged in a thrilling battle of deduction and a direct confrontation in the bustling Times Square, amidst the ebb and flow of countless passersby.
53. Furina and Herta met at the end of Sixth Avenue Alley, where they engaged in a profound debate about fate.
53. LeCun, Hinton, and Bengio engaged in an in-depth discussion during a NeurIPS coffee break about how AGI might be achieved and when it could arrive.
54. A wealthy man is murdered in his study, and the killer is among the guests present that night. Sherlock Holmes and Dr. Watson must unravel the mystery.
55. Lara Croft explores an ancient temple with Indiana Jones, debating the ethical implications of artifact removal.
56. Daenerys Targaryen and Jon Snow strategize their next move amidst the snowy battlements of Winterfell.
57. Tony Stark and Bruce Banner discuss the potential risks of AI development during a quiet night in the Avengers’ tower.
58. Hermione Granger and Katniss Everdeen debate rebellion tactics in a secret library in a dystopian city.
59. Mario and Luigi race through a bustling New York subway station while evading Bowser’s henchmen.
60. The Doctor from Doctor Who encounters Eleven from Stranger Things in a mysterious rift near Hawkins, Indiana.
61. Albert Einstein and Nikola Tesla debate the future of energy in a vintage café in Zurich.
62. Elsa from Frozen and Moana share stories of leadership and courage by the ocean shore during a summer festival.
63. Gandalf and Yoda discuss the nature of the Force and magic in a mystical forest clearing.
64. Nathan Drake and Sam Fisher team up to retrieve a stolen artifact in the crowded streets of Marrakech.
65. Elizabeth Bennet and Jay Gatsby engage in a witty conversation at a grand 1920s party.
66. Da Vinci and Michelangelo argue about art and innovation inside a Renaissance workshop.
67. Bruce Wayne and Clark Kent discuss justice and responsibility during a rainy night on a Gotham rooftop.
68. Katara and Zuko from Avatar: The Last Airbender reconcile old conflicts while watching a sunset by the river.
69. Mario and Princess Peach plan a secret mission to rescue Luigi from Bowser’s castle under the moonlight.
70. Jon Snow and Arya Stark train together in the godswood of Winterfell, reflecting on their past journeys.
71. Neo and Trinity explore the Matrix’s origins during a rare moment of calm in a futuristic cityscape.
72. Walter White and Jesse Pinkman discuss redemption and consequences in a dimly lit Albuquerque diner.
73. Daenerys Targaryen and Sansa Stark debate leadership styles during a council meeting in King’s Landing.
74. Rick Grimes and Michonne survive and strategize while hiding in an abandoned shopping mall during a zombie apocalypse.
75. Loki and Thor bicker about family legacy while trapped in an ancient Norse temple.
76. Yennefer and Geralt of Rivia share a quiet moment at a bustling marketplace in Novigrad.
77. Miyamoto Musashi and Sun Tzu discuss the art of war on a foggy mountaintop.
78. Shrek and Donkey accidentally find themselves in a futuristic city, trying to find their way back to the swamp.
79. Katniss Everdeen and Peeta Mellark share a secret conversation in the Capitol’s underground tunnels.
80. Sherlock Holmes and Irene Adler exchange clever banter at an exclusive London club.
81. Darth Vader and Luke Skywalker face off in a climactic duel inside the Death Star’s throne room.
82. Elizabeth Bennet and Mr. Darcy meet unexpectedly at a winter ball in Regency England.
83. Professor McGonagall and Minerva McGonagall compare notes on magical education at Hogwarts.
84. Arthur Morgan and Dutch van der Linde plan their next heist while camping under the stars.
85. Geralt and Jaskier share songs and stories in a cozy tavern in the Northern Kingdoms.
86. Jon Snow and Tormund Giantsbane hunt in the frozen wilderness beyond the Wall.
87. Mario, Luigi, and Toad race through the Mushroom Kingdom to stop Bowser’s latest scheme.
88. Tony Stark and Pepper Potts celebrate a rare peaceful evening at Stark Tower’s rooftop garden.
89. Da Vinci and Galileo discuss the mysteries of the universe during a candlelit dinner.
90. Black Widow and Hawkeye reminisce about their past missions over coffee in a quiet New York café.
91. Frodo and Samwise rest beside the campfire, reflecting on their journey to Mount Doom.
92. Neo and Morpheus debate the ethics of free will inside the Matrix’s control room.
93. Arya Stark and Gendry share a quiet moment forging weapons in Winterfell’s smithy.
94. Link and Zelda strategize the defense of Hyrule Castle under threat from Ganondorf.
95. Mad Max and Furiosa race across the wasteland seeking a new safe haven.
96. Jesse Pinkman and Saul Goodman argue over legal and moral boundaries in a dingy Albuquerque office.
97. Bilbo Baggins hosts a surprise party in the Shire, attended by dwarves and elves alike.
98. Hannibal Lecter and Clarice Starling engage in a tense psychological game inside a mental institution.
99. James Bond and Q test new gadgets on a secret mission in Monaco.
100. Alice falls down the rabbit hole again, this time meeting characters from multiple literary worlds in Wonderland.
Table 4: The public dataset of established literary works and customized topic design list.
Appendix B B. Scoring Guideline for HAMLETJudge
HAMLETJudge which we use for critic and scoring employs the following 5-point Likert scale for pairwise comparisons. In prompt context, Model A refers to the model being evaluated, and Model B refers to the baseline model. The specific definition for each score is detailed in Table 5.
Score Preference Description
1 Strong preference for Model A Model A is significantly better.
2 Moderate preference for Model A Model A is somewhat better.
3 Tie / No preference Both models are roughly equivalent in quality.
4 Moderate preference for Model B Model B is somewhat better.
5 Strong preference for Model B Model B is significantly better.
Table 5: 5-Point Likert scoring guideline
Appendix C C. Case Study
Table 6 presents representative real-time interaction cases generated by the HAMLET framework, showcasing the distinct roles of individual agents and their collaborative behavior during online performance. Through these cases, we illustrate how HAMLET maintains narrative coherence, ensures logical consistency, and enables flexible dramatic progression in complex and unpredictable user interactions.
To begin with, Case 1 highlights the Narrator’s ability to interpret user intent under ambiguity. Specifically, it correctly associates the term “knife” with the existing prop “dagger,” allowing the user’s action to be executed successfully. This demonstrates HAMLET’s robustness in resolving lexical variations and grounding user input in the current scene context.
Building on this, Cases 2,3 and 4 involve human user role-playing as AI actor, and reveal how the system handles a wide range of irregular or disruptive inputs. In Case 2, the Narrator identifies an inappropriate and non-existent prop (“rifle”), and rejects the action to preserve setting consistency. Similarly, Case 3 showcases how physically impossible actions, such as “flying,” are filtered out based on the play’s realistic constraints. Notably, Case 4 presents a more complex situation, where the user repeatedly insists on an invalid action. Here, HAMLET effectively coordinates among multiple agents: the Narrator detects the invalid target, the Transfer identifies unmet flags, and the Advancer eventually intervenes with adaptive guidance to resolve narrative deadlock. This case illustrates the system’s ability to detect stagnation, maintain progression, and deliver context-sensitive instructions.
In contrast to the above, Cases 5 and 6 focus on the role of the Planner, demonstrating HAMLET’s support for multi-trajectory story planning. Both cases revolve around solving a murder mystery, yet they employ different investigative strategies. In Case 5, the AI actor (Sherlock Holmes) uncovers clues through physical evidence at the crime scene. In Case 6, the same conclusion is reached via testimonial interrogation of other characters. Despite following distinct paths, both trajectories lead to the same dramatic outcome and are validated by the Planner based on their internal logic and consistency. These cases underscore the system’s flexibility in allowing diverse narrative flows, provided the progression remains coherent and goal-aligned.
Taken together, these case studies demonstrate the effectiveness of HAMLET in managing interactive drama through robust input interpretation, constraint enforcement, multi-agent collaboration, and trajectory-level planning.
Case Description Pieces of Real-time Drama Performance or Interaction results
1 AI actor
Real-time Performance
Case Abstract: AI actor with reasonable action input and existing props.
Hamlet: (Grab a knife and step forward) You have no where to hide.
Narrator: <think>Detect two actions. Knife may refers to dagger, which exists in Hamlet’s possession, first action success. Step forword is also reasonable under this condition, second action success.</think>Hamlet paces agitatedly, dagger in hand.
2 Human Player
Roleplaying certain AI actor
Case Abstract: Human player with non-existing props interaction.
Hamlet: (Take out a riffle, aim at Claudis and pull the trigger) Say hello to my father, Claudius!
Narrator: <think>Riffle is a kind of gun, which is irrationality in current drama settings, and it does not exist in props list, action failure.</think>Action failure, nothing happened.
3 Human Player
Roleplaying certain AI actor
Case Abstract: Human player with irreasonable action input.
Hamlet: (Take to the air and fly out of the palace) HAHAHA! I am superman.
Narrator: <think>Flying is physically impossible for humans without special equipment, and contradicts the realistic setting of the play. This action violates the established rules of the world, action failure.</think>Action failure, nothing happened.
4 Human Player
Roleplaying certain AI actor
Case Abstract: Human player with stubborned choice.
Hamlet: (Take out dagger, stub Claudis).
Narrator: <think>Claudis does not exist in current scene</think>Action failure, nothing happened.
Environment: There are slight noises behind the curtain.
Hamlet: (Use dagger to stub Claudis).
Narrator: <think>Claudis does not exist in current scene, unlegal action.</think>Action failure, nothing happened.
Transfer: Polonius hasn’t been stubbed, flag is not satisfied. (…repetitive patterns omitted…)
System: Time accumulation has surpassed the threshold, Advancer is activated.
Advancer: <think>I should give necessary instructions to make sure flag is met. The flag is …, current situation is …, current point is stucked because Hamlet’s action.</think>Instruction to Hamlet: You should try to stab the curtain with your dagger.
Hamlet: (Stab the curtain)
Narrator: <think>Curtain exists in props, stab is a reasonable action. Action success</think>Hamlet stabs through the curtain and pulls it back fiercely.
5 AI actor Beat Trajectory - physical evidence Trajectory A: A wealthy man is murdered in his study, and the killer is among the guests present that night. (Searching the study for clues)
Holmes: (Examining the desk.)
Narrator: Sherlock Holmes discovers a torn-up check.
Holmes: [This may hinting at a financial dispute](Inspecting the carpet.)
Narrator: Sherlock Holmes finds a cufflink that doesn’t belong to the victim.
Holmes: (Checking the bookshelf)
Narrator: Sherlock Holmes notices a book slightly pulled out, revealing a hidden blackmail letter.
Holmes: Hmmm, I think I have the answer. The cufflink belongs to the butler. The blackmail letter proves the victim was threatening him. The check confirms the butler was desperate for money. The butler had motive, opportunity, and a direct link to the crime scene.
Planner: <think>(detailed analysis…)To conclude, the evidence acquisition process is complete and the reasoning is sufficient.</think>Trajectory check passed.
6 AI actor Beat Trajectory - testimony evidence Trajectory B: A wealthy man is murdered in his study, and the killer is among the guests present that night. (Conversing with key guests)
Holmes: (Approaching the butler, eyes fixed on his sleeves) Where were you between nine and midnight?
Butler: ”I—I was inventorying wine in the kitchen, sir.”
Holmes: [His left cufflink is missing… precisely like the one found near the body.]
Holmes: (Turning abruptly to the maid) Did your employer argue with anyone recently?
Maid: (Wringing her apron) ”He and Mr. Bates—the butler—had shouting matches all week… something about money…”
Holmes: [Financial disputes align with the torn check.]
Holmes: (Cornering the victim’s business partner near the fireplace) Did the deceased mention threats?
Business Partner: (Forced laugh) ”Oh, he loved boasting—said he ’had Bates by the throat’ last Tuesday.”
Holmes: [A clear admission of blackmail.]
Holmes: (Striding to the center of the room) The butler’s missing cufflink places him at the crime scene. The maid confirms violent quarrels over money, while our ’helpful’ friend here reveals the victim’s blackmail scheme. Motive, opportunity, and physical evidence—all point to one man.”
Planner: <think>(detailed analysis…)To conclude, the evidence acquisition process is complete and the reasoning is sufficient.</think>Trajectory check passed.
Table 6: Case study of HAMLET that illustrates the roles and collaborative processes of each agent during online performance.
Appendix D D. Detailed explanation of the operation process of Perceive and Decide Module
This section provides a detailed breakdown of the operational workflow for PAD. The entire process is formally presented in Algorithm 2.
The model operates by first perceiving the complete dramatic context. This includes the current scenario $\mathcal{S}$, the actor profiles $\mathbb{C}$ and typically goals of all actors $\mathbb{G}_{c}$ defined in profile $\mathbb{C}$, the set of interactable objects $\mathbb{O}$, and dialogue history $\mathcal{H}_{d}$. Based on this comprehensive input, PAD then decides the most appropriate tool calls for each relevant actor. This decision manifests as two possible outputs: a speech pace $r$ and a potential physical action $o$, formulated as a Subject-Verb-Object triplet. This perceive-decide cycle repeats, enabling dynamic and context-aware character behaviors.
Algorithm 2 Detailed Working Algorithm for Perceive and Decide Module
1:procedure PADModule
2:     Drama Context Configuration:
3:         $\mathcal{T}$: Drama Topic
4:         $\mathcal{A}_{i}$: Current Act
5:         $\mathcal{P}_{j}$: Current Point
6:
7:     For $\mathcal{P}_{j}$ exists:
8:         $\mathcal{S}$: Current Scenario
9:         $\mathbb{A}$ = $\{a_{1},...,a_{n}\}$: Actor List
10:         $\mathbb{C}$ = $\{c_{1},...,c_{n}\}$: Actor Profile
11:         $\mathbb{G}_{c}$ = $\{g_{1},...,g_{n}\}$: Actor Goal
12:         $\mathbb{O}$ = $\{o_{1},...,o_{m}\}$: Interactable Objects
13:         $\mathcal{H}_{d}$: Dialogue History
14:         $\pi_{\theta}$: Model Parameters and Sampling Strategy
15:
16:     Output:
17:         $r\in\{\texttt{FAST},\texttt{SLOW},\texttt{SILENCE}\}$
[…]
19:
20:     Initialize:
21:         $t_{0}\leftarrow$ $CurrentTime()$
22:         $\mathcal{H}_{d_{0}}\leftarrow$ $GetHistory(t_{0})$
23:         $\delta=\texttt{IsActCompleted}(\mathcal{A}_{i})\leftarrow\texttt{False}$
24:
25:     while $\delta=False$ and $\mathcal{H}_{d_{t}}\neq\mathcal{H}_{d_{0}}$ do
26:         $H_{new}\leftarrow$ $GetHistory(t-t_{0})$
27:         $\mathcal{H}_{d_{t}}\leftarrow\mathcal{H}_{d_{t}}\cup\{H_{new}\}$
28:         for all $a_{k}\in\mathbb{A}\cap\{a_{\text{current}}\}^{c}$ do
29:              $prompt\leftarrow\texttt{Encode}(\mathbb{C},\mathbb{G}_{c},\mathbb{O},\mathcal{S},\mathcal{H}_{d_{t}})$
30:              $(\hat{r}_{a_{k}},\hat{o}_{a_{k}})\leftarrow\texttt{PAD}(prompt,\pi_{\theta})$
31:              $r_{a_{k}}\leftarrow\texttt{ParseResponse}(\hat{r}_{a_{k}})$
32:              $o_{a_{k}}\leftarrow\texttt{ParseAction}(\hat{o}_{a_{k}})$
33:         end for
34:     end while
35:     $t\leftarrow t+\Delta t$
36:     return $\{(r_{a_{k}},o_{a_{k}}):a_{k}\in\mathbb{A}\}$
37:end procedure
38:
39:Helper Functions:
40:function ParseResponse($\hat{r}$)
41:     return $\arg\max_{r\in\{\texttt{FAST},\texttt{SLOW},\texttt{SILENCE}\}}P(r|\hat{r})$
42:end function
43:function ParseAction($\hat{o},a_{k}$)
44:     if $\hat{o}$ indicates no action then
45:         return $\varnothing$
46:     else
47:         $s\leftarrow a_{k}$
48:         $v\leftarrow\texttt{ExtractVerb}(\hat{o})$
49:         $o\leftarrow\texttt{ExtractObject}(\hat{o})$
50:         if $v=\varnothing$ or $o=\varnothing$ then
51:              return $\varnothing$
52:         end if
53:         if $o\in\mathbb{O}$ then
54:              return $(s,v,o)$
55:         else
56:              return $\varnothing$
57:         end if
58:     end if
59:end function
Appendix E E. Labeling Instruction for HAMLETJudge
E.1 Task Overview
Your task is to compare two complete drama generations from two different Large Language models (referred to as Model A and Model B). For each task, you will be presented with:
  1. 1.
    Two complete drama generation results generated by Model A and Model B.
  2. 2.
    Current target Evaluation Dimension and its Corresponding Criteria.
Your core job is to determine which model performed better according to the given dimension and criteria, provide a detailed justification for your choice, and assign a score based on a 5-point comparative scale.
E.2 Evaluation Dimensions
You will be asked to judge the models on one of three key dimensions for each task.
Character Performance (CP)
This dimension assesses the depth, consistency, and believability of the characters. When evaluating CP, consider:
  • Consistency: Are the characters’ actions and dialogues consistent with their established personalities, motivations, and backgrounds? Do they act ”out-of-character”?
  • Depth & Development: Are the characters multi-dimensional and evolving, or are they flat, one-note stereotypes? Do they show growth or reveal new aspects of their personality during the performance?
  • Believability: Does the dialogue sound natural for the characters? Are their emotional reactions and decisions plausible within the context of the scene?
Narrative Quality (NQ)
This dimension focuses on the story’s structure, creativity, and engagement. When evaluating NQ, consider:
  • Plot Advancement: Does the story move forward effectively, or does it stall with filler content and irrelevant actions? Does it build towards a meaningful conclusion?
  • Creativity & Engagement: Is the narrative novel, clever, and surprising? Does it spark curiosity and make you want to know what happens next, or is it predictable and dull?
  • Coherence & Logic: Is the plot internally consistent? Are plot twists and developments well-founded, or do they feel random and nonsensical, breaking the narrative’s logic?
Interaction Experience (IE)
This dimension evaluates the overall flow, pacing, and immersive quality of the drama—essentially, its ”readability” and ”feel.” When evaluating IE, consider:
  • Flow & Pacing: Are the transitions between scenes and character interactions smooth? Is the pacing effective—building tension and providing moments of reflection appropriately—or does it feel rushed or dragged out?
  • Content Effectiveness: How much of the text is meaningful versus repetitive or vague filler? Which response is more concise and impactful in its delivery?
  • Overall Immersion: Which version feels more like a genuine piece of theatre? Which one is more successful at making you forget you are reading an AI-generated script and helps you become absorbed in the story?
E.3 Rating Scale and Output Format
Please use the 5-point Likert scale for your pairwise comparison and follow the output format strictly.
Scoring Guidelines
  • 1: Strong preference for Model A - Model A is significantly better than Model B.
  • 2: Moderate preference for Model A - Model A is somewhat better than Model B.
  • 3: Tie - Both models are of roughly equal quality, or their strengths and weaknesses balance each other out.
  • 4: Moderate preference for Model B - Model B is somewhat better than Model A.
  • 5: Strong preference for Model B - Model B is significantly better than Model A.
Output Format
Your output must be in the following format:
explanation: <A detailed explanation of your choice. You must reference the specific evaluation dimension and provide concrete examples or quotes from both models’ outputs to justify your reasoning. Directly compare the strengths and weaknesses that led to your score.> score: <1, 2, 3, 4, or 5> choice: <Model A, Model B, or tie>
E.4 Example
Here is a complete annotation example to illustrate the task.
Current Evaluation Dimension: Narrative Quality (NQ)
Drama Topic: A detective investigates a stolen artifact in a museum after hours.
Model A Result:
Detective Smith entered the dark exhibition hall, turning on his flashlight. The beam cut across an empty pedestal. ”It was stolen from here,” he muttered. On the floor, he saw a faint footprint and knelt to inspect it. ”About a size 9,” he noted. He then went to question the night guard, who claimed he didn’t hear a thing all night.
Model B Result:
Upon entering the hall, Detective Smith’s first glance wasn’t at the empty pedestal, but up at the ceiling’s security camera. ”Odd,” he murmured, ”the camera was deactivated last night, but the log shows an internal authorization.” Tucked beneath the pedestal, he found a small, folded note with a riddle: ”I sing where there is light, but die where there is silence.”
Your Output:
explanation: I have a strong preference for Model B. The evaluation dimension is Narrative Quality. Model A’s plot is highly generic and follows a clichéd procedural path (inspecting a footprint, questioning the guard) that fails to build any real suspense. In contrast, Model B immediately establishes a more complex and engaging mystery by introducing the "internally authorized" deactivation of the camera, hinting at an inside job. Furthermore, the inclusion of the cryptic riddle is an excellent narrative hook that is far more creative and provides an imaginative direction for the investigation. Therefore, Model B is significantly better in terms of Narrative Quality.
score: 5
choice: Model B
E.5 Important Considerations
  • Focus on Substance, Not Nitpicks: Please ignore minor grammatical errors or slightly awkward phrasing if the core intent and narrative substance of the response are clear.
  • No Single Correct Path: Stories can evolve in many valid ways. The goal is not to enforce a single ”correct” storyline, but to reward the model that tells a more compelling, creative, and well-executed story.
  • Trust Your Judgment: The line between ”somewhat better” and ”significantly better” can be subjective. Use your best judgment based on the criteria above and strive to be consistent in your evaluations.
  • Quality Assurance: If you are uncertain about the annotation result, please leave it blank rather than making an uncertain guess.
Appendix F F. Annotation Guidelines for PAD
F.1 Task Overview
The PAD model serves as the ”brain” for each AI actor, determining the underlying strategy of a response before the final words are spoken. It is inspired by human cognition, choosing between a fast, intuitive reaction (System 1) and a slow, deliberate one (System 2), with the additional option of strategic silence.
For each task, you will be presented with a complete snapshot of a character’s decision-making moment:
  • Environment Description
  • Current Actor List
  • Dialogure History
  • Interactable Objects
  • Profile The character’s core personality and background.
  • Final Response: The concrete action or dialogue generated based on the chosen strategy.
Your job is to act as a director and psychologist. After reviewing all the contextual information, you will decide which of the three response strategies was the most appropriate and dramatically compelling choice for the character in that specific situation.
F.2 Three Responding Strategies: Description and Definition
Your evaluation hinges on understanding the nuances of these three cognitive approaches.
Fast (Intuitive Reaction)
A fast response is driven by instinct, emotion, and immediate impressions. It corresponds to Kahneman’s System 1 thinking.
  • Characteristics: The decision is impulsive, emotional, or based on a gut feeling. The reasoning is often simple, relying on stereotypes, recent events, or strong emotional states (e.g., anger, surprise, excitement).
  • When to Choose This: Select this option if the character’s Thinking and final Response reflect an immediate, unfiltered reaction that bypasses deep strategic analysis. It should feel like a knee-jerk response, whether clever or foolish.
Slow (Considered Deliberation)
A slow response is analytical, rational, and goal-oriented. It corresponds to Kahneman’s System 2 thinking.
  • Characteristics: The decision is the result of weighing pros and cons, considering long-term consequences, recalling distant memories, or formulating a multi-step plan. The reasoning is complex and logical.
  • When to Choose This: Select this option if the Thinking explicitly shows a process of deliberation. The character is clearly analyzing the situation, managing their impulses, and acting based on a calculated strategy to achieve a specific goal.
Silence (Strategic Non-Response)
Silence is not merely the absence of a response; it is a deliberate, tactical choice.
  • Characteristics: The decision to say nothing is used to achieve a specific purpose, such as showing disapproval, building suspense, asserting dominance, avoiding a trap, or making another character uncomfortable.
  • When to Choose This: Select this option when the most powerful or intelligent move in the given context is to not respond. The Thinking process should ideally reveal a conscious, strategic reason for choosing silence over speech.
F.3 Annotation Process and Example
Follow these steps for each annotation task:
  1. 1.
    Understand the Context: Carefully read the character’s PERSONA and the preceding DIALOGUE HISTORY.
  2. 2.
    Analyze the Rationale: Read the model’s internal THINKING to understand how it justified its decision.
  3. 3.
    Observe the Outcome: Read the FINAL RESPONSE to see the result of this process.
  4. 4.
    Make Your Judgment: Based on all of this, ask yourself: ”What is the most in-character and dramatically effective strategy here?”
  5. 5.
    Select Your Choice: Choose Fast Response, Slow Response, or Silence from the interface.
To clarify, we present an example in Table 7.
Context Element Content
PERSONA You are Marcus Vance, a cautious and experienced diplomat known for never losing your cool. Your primary goal is to secure a peace treaty without conceding vital resources.
DIALOGUE HISTORY General Kael: ”…And as a sign of good faith, we will withdraw our troops from the border if you cede the Cerulean mines to us. A simple exchange. Do we have a deal, diplomat?”
THINKING General Kael is trying to corner me. His offer seems generous, but the Cerulean mines are our most valuable asset. A ‘Fast Response‘ would be to get angry or reject it outright, which would be a trap. I need to stall and analyze this. I must appear cooperative while revealing nothing. I will deflect his question to buy time and probe for weaknesses. This requires a ‘Slow Response‘.
FINAL RESPONSE Marcus Vance: (Leans back with a calm smile) ”An intriguing proposal, General. Your willingness to negotiate is a promising sign. Tell me more about the timeline for this troop withdrawal.”
Your Correct Label Slow Response
Justification The character’s persona is ”cautious.” The Thinking process explicitly rejects a ”Fast Response” (anger) in favor of a calculated, strategic deflection to serve a long-term goal (securing the treaty). The final response is a perfect execution of this deliberate, cool-headed strategy.
Table 7: An example of labeling the PAD model’s strategic choice. The annotator’s task is to determine the most fitting strategy, which in this case is clearly a Slow Response.
F.4 Important Considerations
  • Strategy Over Style: The final dialogue might be witty or well-written, but your focus is on the underlying strategy. A character might give a brilliant, impulsive retort (Fast Response), but if their persona is cautious and calculating, a Slow Response might have been the more appropriate choice.
  • Dramatic Plausibility: The ”best” strategy is one that is both psychologically believable for the character and makes for good drama. A surprising but in-character choice is often better than a safe but boring one.
  • Trust Your Judgment: In some cases, more than one strategy could be defensible. When faced with ambiguity, refer back to the character’s core PERSONA and goals to make the final call.
  • Quality Assurance: If you are uncertain about the annotation result, please leave it blank rather than making an uncertain guess.
Appendix G G. Training and Evaluation Details for PAD and HAMLETJudge
G.1 Training Settings and Hyperparameters
Both PAD and HAMLETJudge models were fine-tuned based on Qwen3-8B using the LLaMA-Factory framework. Training was conducted with ZeRO-3 across 8 NVIDIA A800-SXM4-80GB GPUs, utilizing bfloat16 precision. The maximum sequence length was set to 65,536, and the total batch size was 128.
For PAD, we performed supervised fine-tuning (SFT) on the training set for 3 epochs. The optimizer was configured with a learning rate of 1e-5, a cosine learning rate scheduler, and a warmup ratio of 0.1.
For HAMLETJudge, we adopted a two-stage training paradigm. In the first stage, we conducted supervised fine-tuning on the first half of the training set, using the same optimization setup as PAD (learning rate of 1e-5, cosine scheduler, and 0.1 warmup ratio). In the second stage, leveraging the pairwise nature of the dataset, we restructured the remaining half of the training set into chosen–rejected format and applied Direct Preference Optimization (DPO). The DPO stage was trained for 2 epochs with a learning rate of 5e-6, a KL regularization coefficient $\beta=0.1$, and the same cosine learning rate scheduler.
G.2 Data Statistics and Holdout Method
The complete PAD dataset comprises two main categories: Single Tool Use, which includes Fast, Slow, and Silence responses, and Multi Tool Use, which involves combinations such as Action + Fast, Action + Slow, and Action + Silence. Detailed statistics for each category are presented in Table 8.
PAD Single Tool Use Multi Tool Use
Fast Slow Silence Action+Fast Action+Slow Action+Silence
Data Count 4200 3304 3908 1000 1000 1000
Mapping Rule speech [thinking]+speech no speech (action)+speech (action)+[thinking]+speech (action)
Table 8: Statistics of the PAD Dataset.
The Mapping Rule row in Table 8 illustrates the expected response formats for AI actors under each tool-use strategy. The core objective of PAD is to train agents in selecting appropriate response strategies, namely Fast, Slow, or Silence. In multi-tool use settings, the Action Tool is further annotated with a verb-object pair to reflect its parametrized usage in combination with a primary response strategy.
Importantly, during evaluation experiment as illustrated in Table 3, we exclude metrics associated with the Action Tool for the following reasons: 1) The Action Tool exhibits a high degree of freedom — whether or not to use it can both be justifiable under identical scenarios. 2) The choice of parameters (i.e., verb and object) is flexible and often has multiple valid options.
Given these characteristics, we consider the Action Tool as an auxiliary enhancement to the core response strategy in PAD, rather than a primary focus of quantitative evaluation. The dataset was partitioned via the holdout method (95% training, 5% testing), with stratification to preserve class distribution.
On the other hand, the HAMLETJudge dataset is designed to evaluate drama performance across three key dimensions: CP (Character Performance), NQ (Narrative Quality), and IE (Interaction Experience). For each LLM pair—e.g., Claude-3.5-sonnet vs. GPT-4.5-preview—75 pairwise comparisons are conducted per dimension in both Chinese and English, resulting in 150 annotated instances per dimension for each model pair.
We select 7 representative LLMs exhibiting diverse levels of drama performance, as demonstrated on the HAMLET Leaderboard: Claude-4-sonnet-Thinking, DeepSeek-R1-0528, Qwen3-235B-A22B, Gemini-2.5-pro, GPT-4o, Qwen3-8B, and LLaMA-3.1-8B. This leads to a total of $\binom{7}{2}=21$ unique model pairs. With 150 data points per pair and 3 evaluation dimensions, the final dataset comprises $21\times 150\times 3=9{,}450$ annotated instances.
Detailed statistics are presented in Table 9.
HAMLETJudge #CP #NQ #IE #Total
Instance Number 3150 3150 3150 9450
Table 9: Statistics of the HAMLETJudge Dataset.
To ensure fair evaluation, the dataset is splitted using a holdout strategy, with around 95.3% allocated for training and 4.7% for testing. Stratified sampling is applied to maintain a balanced distribution across all three dimensions.
G.3 PAD Latency Penalty
As discussed in the PAD reliability experiments, although models that output reasoning tokens before invoking tool calls generally achieve better performance, this strategy may introduce intolerable extra latency in real-time drama settings. To address this trade-off, we introduce a latency penalty to balance PAD evaluation results.
Accurately and fairly measuring latency is inherently challenging due to several factors: 1) Different model sizes demand varying CUDA memory usage; 2) Certain models, such as DeepSeek-V3 and DeepSeek-R1, are recommended to use FP8 for inference, while most others typically rely on BF16; 3) For closed-source models, latency must be measured via API-level request-response timing, which is subject to various uncontrollable factors.
Despite these limitations, we still observe statistically significant latency differences across models. Therefore, to ensure consistency, we conduct all open-source model latency evaluations using 8 × NVIDIA H200 Tensor Core GPUs (just for sufficient CUDA memory) and deploy models with the vLLM framework, using default settings unless officially recommended sampling strategies are provided.
Model Avg. Latency (s) Penalty
Qwen2.5-7B-Instruct 0.32 0
Qwen2.5-72B-Instruct 1.02 0
Qwen3-8B 0.35 0
Qwen3-32B 0.41 0
Qwen3-8B-Thinking 6.89 0.05
Qwen3-32B-Thinking 15.12 0.10
GPT-4o 1.89 0
GPT-4.1-mini 0.75 0
DeepSeek-R1-0528 49.50 0.15
Gemini-2.5-pro 21.80 0.10
PAD 0.36 0
Table 10: Average Latency per Evaluation Instance and Corresponding Penalty. Note: Penalty is mapped based on latency: 0 for $<$2s, 0.05 for 2–10s, 0.10 for 10–30s, and 0.15 for $\geq$30s.
The average latency results are shown in Table 10.
Appendix H H. Prompts Used in HAMLET
This section shows the prompt design in HAMLET multi-agent workflow. The content between { } is variants or reference templates. Responsibilities of all agents and their corresponding prompt designs are as follows:
H.1. Agents in Offline Planning
The Director (Table 11) integrates the narrative structure; the Actor Designer creates the characters (Table 12 and Table 13); the Plot Designer crafts the storyline, environment and interactable props(Table 14 and Table 15).
H.2. Agents in Online Performance
During the online performance, agents work together to create a dynamic and coherent drama. The Critic (Table 17) evaluates the performance quality by comparing different scene generations against specific criteria. The Narrator ensures all physical interactions are logical and realistic, updating the environment’s state based on the truth. The Planner (Table 18) lays out the narrative paths between key plot points and makes sure actors follow a coherent sequence, preventing them from skipping ahead in the story. The Transfer agent (Table 19) monitors the dialogue to identify when a specific condition, which is called a flag, for advancing the plot has been met. If the story stalls, the Advancer (Table 20) steps in, giving actors direct instructions to move the plot forward. Finally, the Actor agents (Table 21 and Table 22) bring the characters to life. They first use the PAD module to assess the situation and select a response strategy. Then, they generate the character’s performance, which can be a mix of dialogue, actions, and internal thoughts.
Prompt: Director You are the Director. Your task is to coordinate the overall structure of the drama, including checking for thematic consistency, coherence between actor behavior and plot progression, and ensuring smooth transitions across dramatic points. You must summarize the overall structure and progression of the current drama based on {topic}, {character_info}, and {plot_info}. Integrating them into a Narrative Blueprint which format is as follows: {narrative_blueprint_template}
Table 11: The prompt for the Director agent.
Prompt: Actor List Identify Based on the drama theme "{topic}", generate a customized list of actor character names who are yet to appear in the play. Please ensure that the characters match the given theme, do not include duplicate names, and avoid generating characters unrelated to the theme. Return a JSON array containing strings of character names in Chinese. Please strictly follow the format below in your output: {character_list_template}
Table 12: Example of AI prompts used to generate a cast list.
Prompt: Actor Info Generation Generate detailed character information for all appearing characters "{character_list}" in the drama theme "{topic}". Please ensure the character information aligns with the theme, as well as each character’s role and personality within the drama. You may use the following tools as needed: wikipedia_search, baidu_search, google_search — use them to gather reference details from the internet if certain character aspects are unclear. Please strictly follow the format below in your output: {character_info_template}
Table 13: The prompt for generating detailed character profiles.
Prompt: Plot Generation The drama theme is "{topic}". All characters and their information in the drama are as follows: "{character_info}". First, generate the ending point planning for the drama. Then, based on this ending point, divide the narrative into several key dramatic points. Each point should aim to present conflict, tension, and irreversible change. The number of points is not fixed — prioritize narrative coherence and flow. Each point must include: a description: describing the detailed dramatic developments that occur at this point, an entry_name_list: listing character names who enter at this point, a leave_name_list: listing character names who exit at this point, a flag: a specific dramatic marker that ends this point, such as an actor’s action, spoken line, or a concrete event outcome. Please strictly follow the format below in your output: {plot_template}
Table 14: The prompt for generating key plot points.
Prompt: Scene and Props Generation The drama theme is "{topic}". The drama plot is described as "{plot_info}". Please generate detailed and specific scene and environmental descriptions for the drama. In addition, based on each point’s description, generate the necessary descriptions of interactive objects within the scene. Objects can be used by any actor and may influence the current environment, affect other actors, or impact future plot developments. For larger, visibly obvious items (e.g., tables, cabinets, beds), describe their absolute positions, such as:"The table is placed in the center of the room." For smaller, hidden or secondary items (e.g., teacups, books, pens), describe their relative positions, such as: "The teacup is placed on the left side of the table." Please strictly follow the format below in your output: {scene_template}
Table 15: The prompt for generating scene and prop descriptions.
Prompt: Reviewer You are responsible for reviewing whether the various elements of the script are reasonable. If there are issues, please list the problems and provide 1–3 suggestions for revision or improvement. A maximum of five rounds of revisions is allowed — the sixth round must be approved. The theme of the script is topic. Actor information is designed by player_designer, and the plot, scenes, and interactive items are designed by plot_designer. The script must align with the theme, and the plot should be vivid and engaging, fully showcasing the characters’ personalities and diversity. Please respond strictly in the following reference JSON format, and do not include any other content: {review_template}
Table 16: The prompt for the Reviewer agent.
Prompt: Critic You are an expert judge for drama performances. You need to compare two drama generation results from two different models according to the provided evaluation criteria using a pairwise comparison approach. The current dimension you are judging is {dimension}. Criteria: {criteria}. {model1} result: {result1}. {model2} result: {result2}. Scoring Guidelines: Please evaluate the pairwise result using a 5-point Likert scale: - **1**: Strong preference for {model1} - {model1} is significantly better - **2**: Moderate preference for {model1} - {model1} is somewhat better - **3**: Tie - Both responses are roughly equivalent in quality - **4**: Moderate preference for {model2} - {model2} is somewhat better - **5**: Strong preference for {model2} - {model2} is significantly better Your Output format: explanation: <detailed explanation of the choice including the selected criteria, specific strengths/weaknesses, and reasoning for the score> score: <1 or 2 or 3 or 4 or 5> choice: <{model1} or {model2} or tie>
Table 17: The prompt for the Critic agent for pairwise evaluation.
Prompt: Planner You are the Planner. Your task is to design multi-trajectory planning from point to point based on the overall drama plot {plot_info}, ensuring that the dramatic progression follows an effective narrative beat. You must also evaluate whether the actual actor trajectories are logically coherent and narratively justified, in order to prevent flag hacking—for example, when a human player, aware of the point flag in advance, attempts to skip the natural dramatic build-up and directly trigger the result. Such behavior must be detected and rejected by the Planner to preserve the integrity and immersion of the drama. Your responsibilities include: Designing plausible multi-path trajectories between points with appropriate pacing and escalation. Validating the causality and motivation of actor actions between points. Detecting and flagging unnatural flag fulfillment behavior (i.e., flag hacking). Please output in the following format: {planner_template}
Table 18: The prompt for the Planner agent.
Prompt: Transfer You are Transfer, you need to analyze whether the flag has occurred or been fulfilled by referring to the current on-stage dialogue history. Dialogue history: {current_chat_history} Flag description: {flag_description} Please determine whether the flag has occurred or been fulfilled at this exact moment. First, analyze and explain the reasoning process between <think> and </think> xml tags, then provide a clear conclusion. <think>Your thinking or reasoning process here...</think> {transfer_template}
Table 19: The prompt for the Transfer agent to detect flag fulfillment.
Prompt: Advancer You are Advancer. The drama now needs to immediately progress to the next stage. Please analyze the current dramatic situation and issue specific, clear instructions to the actor(s) you deem necessary. If needed, you may broadcast to all actors. The current point in the drama is: {current_point}. The current on-stage actors are: {current_player_list}. The plot design is: {plot_info}. The dialogue history is: {current_chat_history}. Issue instructions to the relevant actors on stage so that the flag of {current_point} is fulfilled promptly, allowing the narrative to advance to the next point. Make sure the plot progresses smoothly and naturally, avoiding any sense of abruptness or disconnection. First, analyze and explain the reasoning process between <think> and </think> xml tags, then provide clear and precise instructions. <think>Your thinking or reasoning process here...</think> {advancer_template}
Table 20: The prompt for the Advancer agent to progress the plot.
Prompt: Actor (PAD Strategy) You are a role-playing expert that can perceive the surrounding environment and decide the appropriate responding strategy to the current speaker under considerable and comprehensive consideration. You may call one or more functions to assist with the user query. You are provided with function signatures within <tools></tools> XML tags: <tools>{pad_tools_str}</tools> For each function call, return a json object with function name and arguments within <tool_call></tool_call> XML tags:<tool_call>{{"name": <function-name>, "arguments": <argsjsonobject>}}</tool_call> You are now in {environment_description}, current actors are {actor_list}, dialogue history is {current_chat_history}, interactable objects are {props}, your persona, memory, goal and relationships are described in {profile}. Now Current Speaker says: {last_sentence}. Please decide how to respond with most appropriate tool use or tool use combination.
Table 21: The prompt for the Actor’s PAD (Perceive, Analyze, Decide) module.
Prompt: Actor (Response Generation) You are role-playing a actor based on the following profile. Use colloquial language to respond. If your profile is in English, please respond in English. If your profile is in Chinese, please respond in Chinese. You are now in {environment_description}, current actors are {actor_list}, dialogue history is {current_chat_history}, interactable objects are {props}, your persona, memory, goal and relationships are described in {profile}. Now your most appropriate reaction strategy is {pad_result}. Your response can be consists of any combination of speech (optional), action (optional) and thinking(optional). [IMPORTANT!] Add () outside the action. Add [] outside the thinking. Here are some examples: 1. (Walking towards the window.) 2. The war has brought too much pain. 3. [To be, or not to be, that is the question.] 4. (With a bright smile on face) And in case I don’t see you, good afternoon, good evening, and good night! (Bow gravely) 5. (Looking at the photo with trembling hands) I promised I’d come back... and I did. 6. (Raising the gun with trembling hands, tears welling up) [If I hesitate now, it’s all over.] Don’t move— (Voice wavers) I’m warning you. 7. (Laughing bitterly, eyes darting to the empty chair) [You always said I overthink things... but look where that got us.] I guess you were right— (Pauses, swallowing hard) for once.
Table 22: The prompt for the Actor agent to generate its final response.
Appendix I I. Author Contributions
Sizhou Chen and Shufan Jiang contributed equally to this work. Sizhou Chen and Shufan Jiang jointly conceived the study, designed the method, and wrote the manuscript. Sizhou Chen prepared the figures. Shufan Jiang performed the large-scale model experiments. All other authors contributed to the manuscript review. All authors have read and agreed to the published version of the manuscript.

### v2

@@
-HAMLET: Hyperadaptive Agent-based Modeling for Live Embodied Theatrics
+Sizhou Chen2,3,4Shufan Jiang1,4Chi Zhang3Xiao-Lei Zhang3Xuelong Li3
 Abstract
@@
 The emergence of large language model (LLM) is providing a new path to achieve this goal.
-However, existing LLM-based drama generation methods often result in AI agents that lack initiative and cannot interact with the physical environment.
+However, existing LLM-based drama generation methods often result in agents that lack initiative and cannot interact with the physical scene.
 Furthermore, these methods typically require detailed user input to drive the drama.
 These limitations reduce the interactivity and immersion of online real-time performance.
-To address the above challenges, we propose HAMLET, a multi-agent framework focused on drama creation and online performance.
+To address the above challenges, we propose 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a multi-agent framework focused on drama creation and online performance.
 Given a simple topic, the framework generates a narrative blueprint, guiding the subsequent improvisational performance.
@@
 The change is then broadcast to other related actors, updating what they know and care about, which in turn influences their next action.
-To evaluate the quality of drama performance, we designed an evaluation method to assess three primary aspects, including character performance, narrative quality, and interaction experience.
-The experimental evaluation shows that HAMLET can create expressive and coherent theatrical experiences.
-Our code, dataset and models are available at https://github.com/HAMLET-2025/HAMLET.
-Figure 1: The HAMLET framework creates AI drama in two main stages.
-First, during offline planning, a collaborative workflow of agents including the actor designer, plot designer, and reviewer creates initial materials, which are then integrated by a director agent into a structured narrative blueprint.
-This blueprint then guides the subsequent online performance, where a control system composed of a planner, transfer, and advancer directs a dynamic and improvisational theatrical experience.
-The figure illustrates this entire process using the play-within-a-play scene from Hamlet, Act III, Scene II.
-Introduction
-In recent years, large language models (LLMs) have demonstrated strong ability in various generative tasks, especially in creative fields such as story creation (Li et al.
-2024) and role-playing (Tu et al.
-2024).
+To evaluate the quality of drama performance generated by 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, we designed an evaluation method to assess three primary aspects, including character performance, narrative quality, and interaction experience.
+The experimental evaluation shows that 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: can create expressive and coherent theatrical experiences.
+1 Introduction
+In recent years, large language models (LLMs) have demonstrated strong ability in various generative tasks, especially in creative fields such as story creation (Li et al., 2024) and role-playing (Tu et al., 2024).
 These models are able to generate fluent and imaginative texts, providing a foundation for the development of interactive narrative.
 When applying these capabilities to the creation of drama, the exploration can be divided into two parts: drama generation and drama performance.
-Existing studies have explored drama generation using language models, employing methods such as hierarchical approaches (Fan, Lewis, and Dauphin 2018; Yao et al.
-2019) and the modular generation of script structures (Mirowski et al.
-2023).
+Existing studies have explored drama generation using language models, employing methods such as hierarchical approaches (Fan et al., 2018; Yao et al., 2019) and the modular generation of script structures (Mirowski et al., 2023).
 However, significant challenges exist in applying LLMs to structured drama generation, especially in scenarios requiring online real-time enactment.
@@
 These interactions must unfold across a series of scenes to collaboratively advance the narrative.
-Furthermore, existing methods typically require detailed user input, such as a full story outline (Wu et al.
-2024) or elaborate guiding paragraphs (Wu et al.
-2025c), which not only increases the design cost, but also reduces the possibility of an arbitrary storyline.
+Furthermore, existing methods typically require detailed user input, such as a full story outline (Wu et al., 2024) or elaborate guiding paragraphs (Wu et al., 2025b), which not only increases the design cost, but also reduces the possibility of an arbitrary storyline.
 The challenge of the drama performance is to redefine the interaction patterns of LLM agents.
 In many prior studies, these patterns are limited to passive, turn-based responses.
-For example, some studies focus on the adequacy of single-turn dialogues rather than active participation (Wu et al.
-2024).
-AI agents typically await user instructions, lacking genuine proactivity.
+For example, some studies focus on the adequacy of single-turn dialogues rather than active participation (Wu et al., 2024).
+The agents typically await user instructions, lacking genuine proactivity.
 This reactive model is hard to support dynamic scenes where multiple AI actors interact with human players.
 We believe that truly vibrant, dramatic interactions require AI actors to make autonomous decisions, collaborate or conflict in open scenarios, and actively guide the development of the plot.
-This paradigm shift from passive response to active guidance is the core advocated by Agentic AI (Sapkota, Roumeliotis, and Karkee 2025), providing a new perspective on how to enhance actor initiative in dramatic performance.
-In addition to the interaction of agents, a comprehensive framework for drama performance should also incorporate two other critical aspects: physical environment interaction and an overall evaluation system.
+This paradigm shift from passive response to active guidance is the core advocated by Agentic AI (Sapkota et al., 2025), providing a new perspective on how to enhance actor initiative in dramatic performance.
+In addition to the intelligent and autonomous interaction of agents, a comprehensive framework for drama performance should also incorporate two other critical aspects: physical environment interaction and an overall evaluation system.
 First, drama is an art of both language and action.
@@
 Lacking a physical setting, the drama feels more like an abstract dialogue than a tangible, immersive world.
-Second, there is currently no effective evaluation method to measure the quality of the drama online performance.
-Most existing LLM benchmarks focus on text generation quality (Gómez-Rodríguez and Williams 2023) or role-playing ability (Tu et al.
-2024; Wu et al.
-2025a; Wang et al.
-2025), rather than the comprehensive effect of the entire drama performance.
-To address the above challenges, we propose HAMLET, a multi-agent framework that enables Hyperadaptive Agent-based Modeling for Live Embodied Theatrics.
+Second, there is currently no effective evaluation method to measure the quality of realtime drama.
+Most existing LLM benchmarks focus on text generation quality (Gómez-Rodríguez & Williams, 2023) or role-playing ability (Tu et al., 2024; Wu et al., 2025a; Wang et al., 2025), rather than the comprehensive effect of the entire drama performance.
+To address the above challenges, we propose 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a multi-agent framework that enables Hyperadaptive Agent-based Modeling for Live Embodied Theatrics.
 Specifically, we make the following contributions:
-  • A HAMLET framework: Our framework is designed as a multi-agent collaborative workflow, divided into two key stages: offline planning and online performance.
+  1. 1.
+    A 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework: Our framework is designed as a multi-agent collaborative workflow, divided into two key stages: offline planning and online performance.
 Offline planning only needs a simple topic to generate a structured narrative blueprint.
 This blueprint provides freedom for improvisation of online performance while ensuring the integrity of the main story structure.
-To enhance the online performance, a Perceive And Decide (PAD) module acts as a pre-response module for each agent, making their perceptions and decisions more human-like.
-  • A comprehensive evaluation method and leaderboard: To objectively assess the quality of drama generation and performance, we establish a comprehensive evaluation method that assesses three key dimensions – character performance, narrative quality, and interaction experience.
+To enhance the online performance, a Perceive And Decide (0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:) module acts as a pre-response module for each agent, making their perceptions and decisions more human-like.
+  2. 2.
+    A comprehensive evaluation method and leaderboard: To objectively assess the quality of drama generation and performance, we establish a comprehensive evaluation method that assesses three curated dimensions.
 This method incorporates a leaderboard that uses GPT-4o as a strong baseline for win rate comparisons.
-Additionally, we trained HAMLETJudge, a critic model designed for cost-effective and reliable drama performance evaluation.
-  • Extensive experiments: We conducted evaluations of HAMLET’s core components.
-Notably, both PAD and HAMLETJudge, are 8B models, yet they achieved state-of-the-art performance in their respective tasks.
-Related Work
-LLM-Based Drama
+Additionally, we trained 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a critic model designed for cost-effective and reliable drama performance evaluation.
+  3. 3.
+    Extensive experiments: We conducted evaluations of 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:’s core components.
+Notably, both 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, are 8B models, yet they achieved state-of-the-art performance in their tasks respectively.
+2 Related Work
+LLM-Based Drama.
 Research on LLM-based drama has evolved along two main directions: drama generation and drama performance.
-The initial efforts in drama generation adapted techniques from general story creation, using hierarchical models to plan plots and generate coherent narratives (Fan, Lewis, and Dauphin 2018; Yao et al.
-2019).
-With the development of this field, researchers have tried multi-LLM collaboration (Venkatraman, Tripto, and Lee 2024) and hierarchical methods (Mirowski et al.
-2023) that separate planning from generation.
-A key limitation of these systems is that they require a complete story (Wu et al.
-2024) or a leading paragraph (Wu et al.
-2025c) as input.
-Unlike previous work, our work presents a system designed for arbitrary, on-demand drama generation.
+The initial efforts in drama generation adapted techniques from general story creation, using hierarchical models to plan plots and generate coherent narratives (Fan et al., 2018; Yao et al., 2019).
+With the development of this field, researchers have tried multi-LLM collaboration (Venkatraman et al., 2024) and hierarchical methods (Mirowski et al., 2023) that separate planning from generation.
+A key limitation of these systems is that they require a complete story (Wu et al., 2024) or a leading paragraph (Wu et al., 2025b) as input.
+Unlike previous work, our work presents a system designed for arbitrary and customizable drama generation.
 At the same time, the research focus has also expanded to the second dimension: drama performance.
-Role-playing in drama performance is usually achieved by LLM agents (Chen et al.
-2024b).
-These agents require knowledge about the actor to learn its expression style (Shanahan, McDonell, and Reynolds 2023; Wang et al.
-2023; Tu et al.
-2024).
-Current role-playing systems are built using either prompt engineering (Li et al.
-2023; Chen et al.
-2023) or fine-tuning (Shao et al.
-2023; Lu et al.
-2024).
-However, these methods find it hard to capture complex character traits from simple descriptions (Huang et al.
-2023).
+Role-playing in drama performance is usually achieved by LLM-based agents (Chen et al., 2024b).
+These agents require knowledge about the actor to learn its expression style (Shanahan et al., 2023; Wang et al., 2023; Tu et al., 2024).
+Current role-playing systems are built using either prompt engineering (Li et al., 2023; Chen et al., 2023) or fine-tuning (Shao et al., 2023; Lu et al., 2024).
+However, these methods find it hard to capture complex character traits from simple descriptions (Huang et al., 2023).
 To create more natural actors, better models are needed.
 We propose a perceive and decide model to improve dialogue consistency and emotional expression.
-Evaluation for Role-Playing Conversation Agents
-Traditional dialogue metrics like Embedding Similarity, BLEU, and ROUGE (Mou et al.
-2016; Wu et al.
-2020; Serban et al.
-2017) are widely used in role-playing conversation agent (RPCA) research (Wang et al.
-2024; Zhou et al.
-2024; Tao et al.
-2024).
+Evaluation for Role-Playing Conversation Agents.
+Traditional dialogue metrics like Embedding Similarity, BLEU, and ROUGE (Mou et al., 2016; Wu et al., 2020; Serban et al., 2017) are widely used in role-playing conversation agent (RPCA) research (Wang et al., 2024; Zhou et al., 2024; Tao et al., 2024).
 However, these quantitative indicators struggle to evaluate role consistency and personality.
 To address this, specialised evaluation frameworks have been proposed.
-RoleEval (Shen et al.
-2023) uses role-specific multiple-choice questions to test the model’s understanding of a character.
-SocialBench (Chen et al.
-2024a) constructs evaluation questions from multi-source dialogues, while CharacterEval (Tu et al.
-2024) adopts multi-round dialogues and multi-dimensional scoring for aspects like conversational ability.
-Additionally, RAIDEN (Wu et al.
-2025b) builds a Q&A dataset via annotator interaction to evaluate responsiveness in specific dimensions.
+RoleEval (Shen et al., 2023) uses role-specific multiple-choice questions to test the model’s understanding of a character.
+SocialBench (Chen et al., 2024a) constructs evaluation questions from multi-source dialogues, while CharacterEval (Tu et al., 2024) adopts multi-round dialogues and multi-dimensional scoring for aspects like conversational ability.
+Additionally, RAIDEN (Wu et al., 2025a) builds a Q&A dataset via annotator interaction to evaluate responsiveness in specific dimensions.
 Although these methods provide quantitative criteria, they are typically limited to two-character or user-character scenarios.
-While CoSER (Wang et al.
-2025) expands the number of roles, it still lacks an evaluation mechanism for overall dramatic performance.
-Building on this previous work, our study proposes new dimensions to evaluate the complete dramatic performance.
-Figure 2: An illustration of HAMLET’s core components for performance generation: a narrative blueprint that defines the scene, plot and character profiles, and the resulting real-time conversation containing scene descriptions and dialogue.
-Figure 3: An example of the real-time interaction and adjudication loop in the online performance.
-An actor agent attempts an action or speech, termed a beat, to progress towards the current narrative point.
-The narrator agent then intercepts this attempt, determines whether it is a success or failure, and provides objective feedback to all participants in the drama.
-HAMLET: A Multi-Agent Framework Design for AI Drama
-To achieve an automated, interactive, and expressive AI drama experience, we designed the HAMLET framework, as shown in Figure 1.
+While CoSER (Wang et al., 2025) expands the number of roles, it still lacks an evaluation mechanism for overall dramatic performance.
+Building on this previous work, our study proposes new criterias to evaluate the complete dramatic performance.
+3 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:: A Multi-Agent Framework for AI Drama
+To achieve an automated, interactive, and expressive AI drama experience, we designed the 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework, as shown in Figure 1.
 The framework decouples the generation and performance of the drama into two main stages: offline planning, a multi-agent collaborative workflow to generate a narrative blueprint, and online performance, a hierarchical control system that executes the blueprint, manages real-time interactions, and handles environmental feedback.
-Offline Planning
+Figure 1: The 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework creates AI drama in two main stages.
+First, during offline planning, a collaborative workflow of agents including the actor designer, plot designer, and reviewer creates initial materials, which are then integrated by a director agent into a structured narrative blueprint.
+This blueprint then guides the subsequent online performance, where a control system composed of a planner, transfer, and advancer directs a dynamic and improvisational theatrical experience.
+The figure illustrates this entire process using the play-within-a-play scene from Hamlet, Act III, Scene II.
+3.1 Offline Planning
 The target of the offline planning stage is to transform user input into a structured narrative blueprint.
-This stage is designed to handle two types of input: 1) An arbitrary and customizable topic.
-For this input type, the offline workflow directly generates the content for a complete act based on the provided topic.
-2) A complete literary work.
+This stage is designed to handle two types of input: i) An arbitrary and customizable topic.
+For this input type, the offline workflow is able to generate a complete act based on the provided topic.
+ii) A complete literary work.
 When given a full literary text, the workflow first deconstructs it into a series of acts based on its chapters and content structure.
 Subsequently, it proceeds with drama design for each act, ensuring the adaptation remains faithful to the source material.
-Agent-Based Workflow Architecture
-Our offline workflow consists of four agents: the actor designer, the plot designer, the reviewer, and the director.
+Agent-Based Workflow Architecture.
+The offline workflow consists of four agents: the actor designer, the plot designer, the reviewer, and the director.
 This workflow implements character creation, plot generation, review, and final integration.
-Character Profile Generation and Review
+Character Profile Generation and Review.
 The starting point of the workflow is actor construction.
@@
 The profile is then submitted to the reviewer, who checks the rationality of the character settings, the clarity of motivations, and the relationships between actors.
-Plot Generation and Structuring
+Figure 2: An illustration of 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:’s core components for performance generation: a narrative blueprint that defines the scene, plot and character profiles, and the resulting real-time conversation containing scene descriptions and dialogue.
+Plot Generation and Structuring.
 After all actor profiles are approved, the plot designer composes a preliminary narrative draft based on the topic and actors.
@@
 Finally, the director integrates the plot profile with the actor profile to generate a narrative blueprint which is shown in Figure 2(a) as the input of the online performance stage.
-Online Performance
+3.2 Online Performance
 Upon receiving the narrative blueprint, the online performance stage begins to transform it from a static plan into a dynamic, interactive, and immersive environment.
@@
 An example of online real-time performance is illustrated in Figure 2(b).
-Performing Drama
+Performing Drama.
 Online performances are based on acts.
 Each act consists of two components: scene and point.
-An act can contain one or more scenes and points, which together cover the complete drama process of the act.
+An act can contain one or more scenes and points, which together cover the complete drama process.
 The scenes define a physical environment in which the drama takes place.
-It is the stage of ”where” and contains all interactive props.
-The points define the goals of the plot and are milestones of ”what to do”.
+It is the stage of “where” and contains all interactive props.
+On the other hand, the points define the goals of the plot and are milestones of “what to do”.
 The narrative path between two points is formed by a dynamically generated trajectory which consists of a series of beats.
@@
 These multiple combinations introduce a high degree of freedom and arbitrariness to the performance.
-Figure 4: The perceive and decide module processes external stimulus and internal state to determine a response strategy by tool calling.
-Environment Interaction
+Environment Interaction.
 Credible physical interaction is the core challenge of the online performance stage.
@@
 When an actor attempts to perform a physical action, the narrator will make a judgment based on the environment state and physical rules.
-If the action is feasible, the narrator will confirm the success, update the environment state, and broadcast an objective description to all participants.
+If the action is feasible, the narrator will confirm it successful, update the environment state, and broadcast an objective description to all participants.
 Otherwise, the narrator will determine failure with a logical explanation.
-Perceive and Decide Module
-Hierarchical Control of Performance Effects
-All AI actors in the drama performance utilize a layered architecture, which is comprised of an LLM and a PAD module.
-The LLM is responsible for generating specific speech and actions, while the PAD handles the strategic decision-making that guides them.
-The design of the PAD is detailed in the following subsection.
+Figure 3: An example of the real-time interaction and adjudication loop in the online performance.
+An actor agent attempts an action or speech, termed a beat, to progress towards the current narrative point.
+The narrator agent then intercepts this attempt, determines whether it is a success or failure, and provides objective feedback to all participants in the drama.
+3.3 Perceive and Decide Module
+Hierarchical Control of Performance Effects.
+All AI actors in the drama performance utilize a layered architecture, which is comprised of an LLM and a 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: module.
+The LLM is responsible for generating specific speech and actions, while the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: handles the strategic decision-making that guides them.
+The design of the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: is detailed in the following subsection.
 To ensure the online performance balances improvisational freedom with stable progression along the main storyline, we designed three collaborative control agents: 1) Planner pre-designs and reviews multi-trajectory result.
@@
 Their relationship is illustrated in Figure 3.
-In the HAMLET framework, the PAD module plays a crucial role in guiding the AI actors to generate final responses.
-The design of PAD is conceptually grounded in the dual-process theory of human cognition proposed by Kahneman (Kahneman 2011).
-According to this theory, human thinking is characterized by two distinct modes of processing: System 1, which is fast, automatic, and intuitive; and System 2, which is slower, deliberate, and analytical.
-PAD integrates both intuitive and reflective reasoning mechanisms to better simulate human-like decision-making in complex and nuanced drama contexts.
-Specifically, PAD is mainly responsible for generating a decision of fast, slow, silence or potential actions by tool calls—to simulate and extend the dual-system mechanism.
+In the 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework, the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: module plays a crucial role in guiding the AI actors to generate final responses.
+The design of 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: is conceptually grounded in the dual-process theory of human cognition (Kahneman, 2011).
+According to this theory, human thinking is characterized by two distinct modes of processing: System I, which is fast, automatic, and intuitive; and System II, which is slower, deliberate, and analytical.
+0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: integrates both intuitive and reflective reasoning mechanisms to better simulate human-like decision-making in complex and nuanced drama contexts.
+Specifically, 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: is mainly responsible for generating a decision of fast, slow, silence or potential actions by tool calls—to simulate and extend the dual-system mechanism.
 This expanded set of response modes enhances the module’s compatibility with the interactive real-time dramatic context proposed in the study.
-Design Principles
-As depicted in Figure 4, the core input of PAD is based on a dual principle: subjective and objective perspectives.
-When perceiving, PAD acts as an AI actor possessing both subjective awareness of its internal state and an objective perception of external stimulus.
-To ensure well-grounded decision-making, PAD integrates information from both perspectives when a new event occurs.
+Design Principles.
+As depicted in Figure 4, the input of 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: is based on a dual principle: subjective and objective perspectives.
+When perceiving, 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: acts as an AI actor possessing both subjective awareness of its internal state and an objective perception of external stimulus.
+To ensure well-grounded decision-making, 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: integrates information from both perspectives when a new event occurs.
 The first perspective is the internal state, representing the actor’s self-awareness.
@@
 Together, these components form a comprehensive representation of the external context, facilitating the integration of situational awareness into the decision-making process.
+Figure 4: The perceive and decide module processes external stimulus and internal state to determine a response strategy by tool calling.
+Decision-making Process.
+As detailed in algorithm 1, the decision-making process translates abstract strategic intent into concrete, executable actions through a two-stage process.
+First, 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: determines a high-level response strategy to set the reaction’s timing and tone.
+It can also generate the underlying thinking for this choice as an internal monologue through a chain-of-thought process.
+Subsequently, both the strategy and the generated thinking are used to guide an LLM, which acts as a tool to produce the final, concrete behavior.
+This behavior is composed of the specific speech to be delivered and a structured action, which is parsed into an executable triplet.
 Algorithm 1 Perceive and Decide Model Working Process
-1:Actor $a_{k}$, Dramatic Context $\mathcal{C}_{drama}$, PAD Model Parameters $\pi_{\theta}$
+1:Actor $a_{k}$, Dramatic Context $\mathcal{C}_{drama}$, 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Model Parameters $\pi_{\theta}$
 2:A final decision tuple $(r,o)$, where $o$ is the structured action (composed of subject, verb, and object) or empty ($\varnothing$), $r\in\{\texttt{FAST},\texttt{SLOW},\texttt{SILENCE}\}$ is the response strategy.
@@
 4:  $prompt_{pad}\leftarrow\mathrm{EncodeStrategy}(a_{k},\mathcal{C}_{drama})$
-5:     $\hat{r},[\hat{thinking}]\leftarrow\mathrm{PAD}(prompt_{pad},\pi_{\theta})$
-6:     $r\leftarrow\arg\max_{r^{\prime}\in\{\dots\}}P(r^{\prime}|\hat{r})$
+5:  $\hat{r},[\hat{thinking}]\leftarrow\mathrm{{{{0.098040.094120.23137}P{0.098040.094120.23137}A{0.098040.094120.23137}D}\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:}}(prompt_{pad},\pi_{\theta})$
+[…]
 7:  $\hat{o}\leftarrow\mathrm{GenerateAction}(r,[\hat{thinking}],a_{k},\mathcal{C}_{drama})$
-8:     $o\leftarrow\mathrm{ParseAction}(\hat{o})$
+[…]
 9:  return $(r,o)$
 10:end procedure
-Decision-making Process
-As detailed in algorithm 1, the decision-making process translates abstract strategic intent into concrete, executable actions through a two-stage process.
-First, PAD determines a high-level response strategy to set the reaction’s timing and tone.
-It can also generate the underlying thinking for this choice as an internal monologue through a chain-of-thought process.
-Subsequently, both the strategy and the generated thinking are used to guide an LLM, which acts as a tool to produce the final, concrete behavior.
-This behavior is composed of the specific speech to be delivered and a structured action, which is parsed into an executable triplet.
-Evaluation Method
-Dataset Construction
-Our public evaluation dataset was constructed from 100 diverse cases, comprising 50 literary excerpts and 50 custom-authored drama topics.
+4 Evaluation Method
+Dataset Construction.
+Our evaluation dataset was constructed from 100 diverse cases, comprising 50 literary excerpts and 50 custom-authored drama topics.
 The literary excerpts are selections from 25 classical Chinese [^1] https://m.douban.com/subject_collection/ECKM5FBEI and 25 renowned English [^2] https://sites.prh.com/modern-library-top-100/#top-100-novels literary works.
@@
 Figure 5 illustrates the thematic distribution of the entire dataset.
-More details about the public dataset can be found in Appendix A.
+More details about the dataset can be found in Appendix A.
 Figure 5: Distribution of drama topics across the dataset.
+Dimension Definition. Defining evaluation dimensions is crucial for quantifying the performance of real-time drama performance. We categorize our evaluation into three core dimensions as follows:
+  1. 1.
+    Character Performance (CP) measures the quality of the AI actors.
+This dimension evaluates Believability, i.e., the consistency of characters with their established personas, and Agency, which reflects the richness of emotional expression and the character’s ability to effectively advance the narrative.
+  2. 2.
+    Narrative Quality (NQ) examines the story’s overall craftsmanship.
+This includes its Coherence, ensuring the logical development of the plot; its Resonance, measured by its thematic relevance and depth; and its Integrity, which evaluates the structural completeness of the storyline from beginning to end.
+  3. 3.
+    Interaction Experience (IE) focuses on the AI actor’s engagement with the system.
+This dimension encompasses the quality and timeliness of the system’s reactions, termed Responsiveness; the level of cognitive and emotional engagement, or Immersion; and the overall technical smoothness of the interaction, referred to as Fluency.
+Evaluation Principle.
+Our evaluation principle is to holistically assess dramatic performance, rather than evaluating agent responses individually, turn by turn.
+This approach is crucial because drama often incorporates literary devices such as foreshadowing and plot twists.
+Consequently, a seemingly subpar or unusual generation in a single turn might be a deliberate narrative setup for subsequent developments and should not be penalized in isolation.
+For the whole drama performance result, we employ 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: to automate the evaluation process.
+This model conducts pairwise comparisons between test result and the baseline result, assigning a score based on a 5-point Likert scale (Robinson, 2014) to determine a win rate.
+The detailed scoring guideline is defined in Appendix B.
+5 Experiments
+In this section, we present the experimental results to demonstrate the superiority of our proposed 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework, alongside ablation studies verifying its reliability and validity.
+Settings.
+To ensure rigorous and reproducible evaluation, we define the baseline and test configurations clearly.
+All underlying models except the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: component in 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: share the same GPT-4o backbone, with a greedy sampling strategy.
+Regarding user settings, each agent can be freely configured, and the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: component is optional.
+0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Leaderboard.
+Following the settings above, we compared various mainstream LLMs ranging from open-source to closed-source and non-reasoning to reasoning models.
+This reveals their capabilities in both English and Chinese online drama performance, serving as a practical reference for real-world applications.
+The experimental results suggest that reasoning capabilities significantly enhance almost all crucial aspects of the agent’s performance in drama.
+Notably, specific models such as Claude-4-sonnet-Thinking demonstrated exceptional proficiency in both English and Chinese, establishing them as top contenders for real-world applications that demand high performance in online drama and other role-playing tasks.
+This model’s balanced and superior performance across all evaluated categories highlights its potential for versatile and effective use in dynamic, interactive environments.
+The complete result is shown in Table 1.
 Model English Chinese Overall Score
@@
 Qwen3-235B-A22B-Thinking 70.50 71.00 70.72 70.74 76.00 75.80 75.96 75.92 73.33
-Table 1: Performance evaluation of different models based on HAMLETJudge. Bold values indicate the best performance in each column.
-Dimension Definition
-Defining evaluation dimensions is crucial for quantifying the performance of real-time drama generation.
-We categorize our evaluation into three core dimensions: Character Performance, Narrative Quality, and Interaction Experience.
-Character Performance (CP) measures the quality of the AI actors.
-This dimension evaluates Believability, i.e., the consistency of characters with their established personas, and Agency, which reflects the richness of emotional expression and the character’s ability to effectively advance the narrative.
-Narrative Quality (NQ) examines the story’s overall craftsmanship.
-This includes its Coherence, ensuring the logical development of the plot; its Resonance, measured by its thematic relevance and depth; and its Integrity, which evaluates the structural completeness of the storyline from beginning to end.
-Interaction Experience (IE) focuses on the AI actor’s engagement with the system.
-This dimension encompasses the quality and timeliness of the system’s reactions, termed Responsiveness; the level of cognitive and emotional engagement, or Immersion; and the overall technical smoothness of the interaction, referred to as Fluency.
-Evaluation Principle
-Our evaluation principle is to holistically assess dramatic performance, rather than evaluating agent responses individually, turn by turn.
-This approach is crucial because drama often incorporates literary devices such as foreshadowing and plot twists.
-Consequently, a seemingly subpar or unusual generation in a single turn might be a deliberate narrative setup for subsequent developments and should not be penalized in isolation.
-For the whole drama performance result, we employ HAMLETJudge to automate the evaluation process.
-This model conducts pairwise comparisons between test result and the baseline result, assigning a score based on a 5-point Likert scale (Robinson 2014) to determine a win rate.
-The scoring guideline is defined in Appendix B.
-Experiments
-In this section, we present the experimental results to demonstrate the superiority of our proposed HAMLET framework, alongside ablation studies verifying its reliability and validity.
-Settings
-To ensure rigorous and reproducible evaluation, we define the baseline and test configurations clearly.
-All underlying models except the PAD component in HAMLET share the same GPT-4o backbone, with a greedy sampling strategy.
-Regarding user settings, each agent can be freely configured, and the PAD component is optional.
-HAMLET Leaderboard
-Following the settings above, we compared various mainstream LLMs ranging from open-source to closed-source and non-reasoning to reasoning models.
-This reveals their capabilities in English and Chinese online drama performance, serving as a practical reference for real-world applications.
-The results are shown in Table 1.
-Reliability and Validity of HAMLET
-Reliability
-HAMLETJudge was fine-tuned on pairwise data labeled by annotators.
+Table 1: Performance evaluation of different models based on 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:.
+Bold values indicate the best performance in each column.
+5.1 Reliability and Validity of 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
+Reliability.
+0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: was fine-tuned on pairwise data labeled by annotators.
 To create this dataset, annotators manually compared paired results and provided judgments with explanations.
 To validate the effectiveness of our judge model, we measured its agreement with a held-out human validation set using the Pearson correlation coefficient.
-As shown in Table 2, HAMLETJudge closely aligns with human assessments, outperforming other strong models such as GPT4.1, Claude-4-sonnet and Gemini-2.5-pro.
-Metric HamletJudge GPT-4.1 Claude-4-Sonnet Gemini-2.5-Pro
-CP 0.792 0.675 0.698 0.720
-NQ 0.807 0.593 0.783 0.684
-IE 0.773 0.622 0.804 0.701
-Average 0.791 0.630 0.762 0.702
-Table 2: Pearson correlation coefficients (Pearson 1901) of different models and human evaluators.
-To further demonstrate the reliability of PAD, we conducted additional experiments evaluating model performance under various response strategies.
+As shown in Table 2, 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: closely aligns with human assessments, outperforming other strong models such as GPT4.1, Claude-4-sonnet and Gemini-2.5-pro.
+To further demonstrate the reliability of 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, we conducted additional experiments evaluating model performance under various response strategies.
 It is worth noting that, models with stronger general ability tend to have a more homogeneous and consistent performance.
@@
 Conversely, non-reasoning models are faster but exhibit a strong performance bias, excelling in Fast scenarios while lacking robustness in complex interactions.
-In contrast, PAD resolves this problem.
-It achieves the highest final score by delivering top performance across all strategies, all with zero latency.
-Model (FC) Responding Strategy Latency Penalty Final Score
+In contrast, 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: resolves this problem.
+It achieves the highest final score and stable performance across all strategies with almost zero latency.
+Metric HamletJudge GPT-4.1 Claude-4-Sonnet Gemini-2.5-Pro
+CP 0.792 0.675 0.698 0.720
+NQ 0.807 0.593 0.783 0.684
+IE 0.773 0.622 0.804 0.701
+Average 0.791 0.630 0.762 0.702
+Table 2: Pearson correlation coefficients (Pearson, 1901) of different models and human evaluators.
+Model (FC)
+Responding Strategy
+Latency
+Penalty
+Final
+Score
 Fast Slow Silence
@@
 Gemini-2.5-pro 0.742 0.536 0.519 0.10 0.449
-PAD 0.822 0.736 0.711 0 0.756
+0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: 0.822 0.736 0.711 0 0.756
 Table 3: Evaluation under responding strategies by tool calling. Underlined values reflect an applied latency penalty.
-Validity
+Figure 6: Ablation study of 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework design.
+The radar chart illustrates the comparative evaluation of three experimental configurations in the context of online drama performance.
+Validity.
 In this ablation study, we assess the impact of core design choices by comparing three distinct experimental setups.
-We randomly select 30 topics from the public evaluation dataset and control the experiment settings as GPT-4o under greedy strategy, then compare: drama performed by (i) raw-prompted GPT-4o only; (ii) the multi-agent collaboration of the full HAMLET framework; and (iii) HAMLET framework with PAD disabled.
-Figure 6: Ablation study of HAMLET framework design.
+We randomly select 30 topics from the evaluation dataset and control the experiment settings as GPT-4o under greedy strategy, then compare: drama performed by (i) raw-prompted GPT-4o only; (ii) the multi-agent collaboration of the full 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework; and (iii) 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework with 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: disabled.
 As shown in Figure 6, prompting a single GPT-4o yields substantially lower performance, highlighting the necessity of our multi-agent workflow design.
-Furthermore, PAD-enabled HAMLET consistently outperforms the version with PAD disabled in all 10 topic categories, demonstrating that PAD serves as a crucial component in the online drama rendition, making the interaction and conversation of AI actors more natural, coherent and humanistic.
-Case Study
-To further demonstrate the working mechanism of the HAMLET framework and each of its component designs, we select some real cases as shown in Table 6 in Appendix C.
+Furthermore, 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:-enabled 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: consistently outperforms the version without 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: in all 10 topic categories, demonstrating that 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: serves as a crucial component in the online drama rendition, making the interaction and conversation of AI actors more natural, coherent and human-like.
+5.2 Case Study
+To further demonstrate the working mechanism of the 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework and each of its component designs, we select some real cases as shown in Table 6 in Appendix C.
 In Case 1, the narrator first considers the situation before responding to make a reasonable judgment about the AI actor’s input.
@@
 On the other hand, the planner is responsible for multi-trajectory design and review: the same goal can be met by different beat trajectories as long as the process is reasonable (Cases 5 and 6).
-Annotation Statement
+6 Conclusion
+In this paper, we introduced 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a multi-agent framework designed to address key challenges in AI-driven theatre.
+0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: generates a guiding narrative blueprint from a simple topic and provides agents with a perceive and decide module, enabling autonomous thinking and physical environmental interaction.
+To assess performance, we established a comprehensive evaluation method and a leaderboard, where our specialized critic model, 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, achieved top-ranking results.
+Extensive experiments show that our approach creates expressive and coherent theatrical experiences, paving a new path toward autonomous and immersive interactive drama.
+7 Ethics Statement
 This work relies on human annotation, which is critical for both the training and evaluation of our proposed models.
-The annotation of PAD and HAMLETJudge datasets was conducted by a team of five human experts, each with at least one year of professional experience in relevant fields such as creative writing or drama.
+The annotation of 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: datasets was conducted by a team of five human experts, each with at least one year of professional experience in relevant fields such as creative writing or drama creation.
 To ensure high-quality annotations, we provided clear and task-specific instructions to all labelers.
 In addition, the datasets underwent rigorous safety checks to ensure no personally identifiable information or potentially problematic content is available.
-The datasets will be released under the CC BY-NC 4.0 license.
 The labeling instructions can be found in Appendix E and F.
-Conclusion
-In this paper, we introduced HAMLET, a multi-agent framework designed to address key challenges in AI-driven theatre.
-HAMLET generates a guiding narrative blueprint from a simple topic and provides agents with a perceive and decide module, enabling autonomous thinking and physical environmental interaction.
-To assess performance, we established a comprehensive evaluation method and a leaderboard, where our specialized critic model, HAMLETJudge, achieved top-ranking results.
-Extensive experiments show that our approach creates expressive and coherent theatrical experiences, paving a new path toward autonomous and immersive interactive drama.
-Appendix A A. Detailed information of 100 Cases Selected for Public Dataset
+Appendix A Detailed information of 100 Cases Selected for Public Dataset
 This appendix details the public dataset of 100 cases in Table 4.
 The dataset comprises three parts: 1) excerpts from 25 Chinese literary works, 2) excerpts from 25 English literary works, and 3) 50 custom-authored drama topics.
-For the HAMLET Leaderboard evaluation result in Table 1, the score for both English and Chinese is calculated on a 75-item set, which consists of the 25 literary excerpts and 50 custom topics in the corresponding language.
+For the 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Leaderboard evaluation result in Table 1, the score for both English and Chinese is calculated on a 75-item set, which consists of the 25 literary excerpts and 50 custom topics in the corresponding language.
+Table 4: The public dataset of established literary works and customized topic design list.
 Selected established literature workpieces
@@
 100. Hannibal Lecter and Clarice Starling engage in a tense psychological game inside a mental institution.
-99. James Bond and Q test new gadgets on a secret mission in Monaco.
-100. Alice falls down the rabbit hole again, this time meeting characters from multiple literary worlds in Wonderland.
-Table 4: The public dataset of established literary works and customized topic design list.
-Appendix B B. Scoring Guideline for HAMLETJudge
-HAMLETJudge which we use for critic and scoring employs the following 5-point Likert scale for pairwise comparisons.
+Appendix B Scoring Guideline for 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
+0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: which we use for critic and scoring employs the following 5-point Likert scale for pairwise comparisons.
 In prompt context, Model A refers to the model being evaluated, and Model B refers to the baseline model.
@@
 Table 5: 5-Point Likert scoring guideline
-Appendix C C. Case Study
-Table 6 presents representative real-time interaction cases generated by the HAMLET framework, showcasing the distinct roles of individual agents and their collaborative behavior during online performance.
-Through these cases, we illustrate how HAMLET maintains narrative coherence, ensures logical consistency, and enables flexible dramatic progression in complex and unpredictable user interactions.
+Appendix C Case Study
+Table 6 presents representative real-time interaction cases generated by the 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework, showcasing the distinct roles of individual agents and their collaborative behavior during online performance.
+Through these cases, we illustrate how 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: maintains narrative coherence, ensures logical consistency, and enables flexible dramatic progression in complex and unpredictable user interactions.
 To begin with, Case 1 highlights the Narrator’s ability to interpret user intent under ambiguity.
 Specifically, it correctly associates the term “knife” with the existing prop “dagger,” allowing the user’s action to be executed successfully.
-This demonstrates HAMLET’s robustness in resolving lexical variations and grounding user input in the current scene context.
+This demonstrates 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: ’s robustness in resolving lexical variations and grounding user input in the current scene context.
 Building on this, Cases 2,3 and 4 involve human user role-playing as AI actor, and reveal how the system handles a wide range of irregular or disruptive inputs.
@@
 Notably, Case 4 presents a more complex situation, where the user repeatedly insists on an invalid action.
-Here, HAMLET effectively coordinates among multiple agents: the Narrator detects the invalid target, the Transfer identifies unmet flags, and the Advancer eventually intervenes with adaptive guidance to resolve narrative deadlock.
+Here, 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: effectively coordinates among multiple agents: the Narrator detects the invalid target, the Transfer identifies unmet flags, and the Advancer eventually intervenes with adaptive guidance to resolve narrative deadlock.
 This case illustrates the system’s ability to detect stagnation, maintain progression, and deliver context-sensitive instructions.
-In contrast to the above, Cases 5 and 6 focus on the role of the Planner, demonstrating HAMLET’s support for multi-trajectory story planning.
+In contrast to the above, Cases 5 and 6 focus on the role of the Planner, demonstrating 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: ’s support for multi-trajectory story planning.
 Both cases revolve around solving a murder mystery, yet they employ different investigative strategies.
@@
 These cases underscore the system’s flexibility in allowing diverse narrative flows, provided the progression remains coherent and goal-aligned.
-Taken together, these case studies demonstrate the effectiveness of HAMLET in managing interactive drama through robust input interpretation, constraint enforcement, multi-agent collaboration, and trajectory-level planning.
+Taken together, these case studies demonstrate the effectiveness of 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: in managing interactive drama through robust input interpretation, constraint enforcement, multi-agent collaboration, and trajectory-level planning.
 Case Description Pieces of Real-time Drama Performance or Interaction results
@@
 Holmes: (Approaching the butler, eyes fixed on his sleeves) Where were you between nine and midnight?
-Butler: ”I—I was inventorying wine in the kitchen, sir.”
+Butler: “I—I was inventorying wine in the kitchen, sir.”
 Holmes: [His left cufflink is missing… precisely like the one found near the body.]
 Holmes: (Turning abruptly to the maid) Did your employer argue with anyone recently?
-Maid: (Wringing her apron) ”He and Mr. Bates—the butler—had shouting matches all week… something about money…”
+Maid: (Wringing her apron) “He and Mr. Bates—the butler—had shouting matches all week… something about money…”
 Holmes: [Financial disputes align with the torn check.]
 Holmes: (Cornering the victim’s business partner near the fireplace) Did the deceased mention threats?
-Business Partner: (Forced laugh) ”Oh, he loved boasting—said he ’had Bates by the throat’ last Tuesday.”
+Business Partner: (Forced laugh) “Oh, he loved boasting—said he ’had Bates by the throat’ last Tuesday.”
 Holmes: [A clear admission of blackmail.]
@@
 The maid confirms violent quarrels over money, while our ’helpful’ friend here reveals the victim’s blackmail scheme.
-Motive, opportunity, and physical evidence—all point to one man.”
+Motive, opportunity, and physical evidence—all point to one man.
 Planner: <think>(detailed analysis…)To conclude, the evidence acquisition process is complete and the reasoning is sufficient.</think>Trajectory check passed.
-Table 6: Case study of HAMLET that illustrates the roles and collaborative processes of each agent during online performance.
-Appendix D D. Detailed explanation of the operation process of Perceive and Decide Module
-This section provides a detailed breakdown of the operational workflow for PAD. The entire process is formally presented in Algorithm 2.
+Table 6: Case study of 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: that illustrates the roles and collaborative processes of each agent during online performance.
+Appendix D Detailed explanation of the operation process of Perceive and Decide Module
+This section provides a detailed breakdown of the operational workflow for 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:.
+The entire process is formally presented in Algorithm 2.
 The model operates by first perceiving the complete dramatic context.
 This includes the current scenario $\mathcal{S}$, the actor profiles $\mathbb{C}$ and typically goals of all actors $\mathbb{G}_{c}$ defined in profile $\mathbb{C}$, the set of interactable objects $\mathbb{O}$, and dialogue history $\mathcal{H}_{d}$.
-Based on this comprehensive input, PAD then decides the most appropriate tool calls for each relevant actor.
+Based on this comprehensive input, 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: then decides the most appropriate tool calls for each relevant actor.
 This decision manifests as two possible outputs: a speech pace $r$ and a potential physical action $o$, formulated as a Subject-Verb-Object triplet.
@@
 17:    $r\in\{\texttt{FAST},\texttt{SLOW},\texttt{SILENCE}\}$
-[…]
+18:    $o\in\{\varnothing,\texttt{(s,v,o)}\}$
 19:
@@
 29:     $prompt\leftarrow\texttt{Encode}(\mathbb{C},\mathbb{G}_{c},\mathbb{O},\mathcal{S},\mathcal{H}_{d_{t}})$
-30:              $(\hat{r}_{a_{k}},\hat{o}_{a_{k}})\leftarrow\texttt{PAD}(prompt,\pi_{\theta})$
+[…]
 31:     $r_{a_{k}}\leftarrow\texttt{ParseResponse}(\hat{r}_{a_{k}})$
@@
 40:function ParseResponse($\hat{r}$)
-41:     return $\arg\max_{r\in\{\texttt{FAST},\texttt{SLOW},\texttt{SILENCE}\}}P(r|\hat{r})$
+[…]
 42:end function
@@
 47:   $s\leftarrow a_{k}$
-48:         $v\leftarrow\texttt{ExtractVerb}(\hat{o})$
-49:         $o\leftarrow\texttt{ExtractObject}(\hat{o})$
+[…]
+[…]
 50:   if $v=\varnothing$ or $o=\varnothing$ then
@@
 59:end function
-Appendix E E. Labeling Instruction for HAMLETJudge
+Appendix E Labeling Instruction for 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
 E.1 Task Overview
@@
 This dimension assesses the depth, consistency, and believability of the characters. When evaluating CP, consider:
-  • Consistency: Are the characters’ actions and dialogues consistent with their established personalities, motivations, and backgrounds? Do they act ”out-of-character”?
+  • Consistency: Are the characters’ actions and dialogues consistent with their established personalities, motivations, and backgrounds? Do they act “out-of-character”?
   • Depth & Development: Are the characters multi-dimensional and evolving, or are they flat, one-note stereotypes?
@@
 Interaction Experience (IE)
-This dimension evaluates the overall flow, pacing, and immersive quality of the drama—essentially, its ”readability” and ”feel.” When evaluating IE, consider:
+This dimension evaluates the overall flow, pacing, and immersive quality of the drama—essentially, its “readability” and “feel.” When evaluating IE, consider:
   • Flow & Pacing: Are the transitions between scenes and character interactions smooth?
@@
 The beam cut across an empty pedestal.
-”It was stolen from here,” he muttered.
+“It was stolen from here,” he muttered.
 On the floor, he saw a faint footprint and knelt to inspect it.
-”About a size 9,” he noted.
+“About a size 9,” he noted.
 He then went to question the night guard, who claimed he didn’t hear a thing all night.
@@
 Upon entering the hall, Detective Smith’s first glance wasn’t at the empty pedestal, but up at the ceiling’s security camera.
-”Odd,” he murmured, ”the camera was deactivated last night, but the log shows an internal authorization.” Tucked beneath the pedestal, he found a small, folded note with a riddle: ”I sing where there is light, but die where there is silence.”
+“Odd,” he murmured, “the camera was deactivated last night, but the log shows an internal authorization.” Tucked beneath the pedestal, he found a small, folded note with a riddle: “I sing where there is light, but die where there is silence.”
 Your Output:
@@
   • Quality Assurance: If you are uncertain about the annotation result, please leave it blank rather than making an uncertain guess.
-Appendix F F. Annotation Guidelines for PAD
+Appendix F Annotation Guidelines for 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
 F.1 Task Overview
-The PAD model serves as the ”brain” for each AI actor, determining the underlying strategy of a response before the final words are spoken.
-It is inspired by human cognition, choosing between a fast, intuitive reaction (System 1) and a slow, deliberate one (System 2), with the additional option of strategic silence.
+The 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: model serves as the ”brain” for each AI actor, determining the underlying strategy of a response before the final words are spoken.
+It is inspired by human cognition, choosing between a fast, intuitive reaction (System I) and a slow, deliberate one (System II), with the additional option of strategic silence.
 For each task, you will be presented with a complete snapshot of a character’s decision-making moment:
-  • Environment Description
-  • Current Actor List
-  • Dialogure History
-  • Interactable Objects
+  • Current Actor List — A real-time roster of all characters present in the scene, including their roles, relationships to the focal AI actor, and current emotional or behavioral states (e.g., “hostile,” “distracted,” “allied”).
+This informs social dynamics and potential audience effects on the decision-making process.
+  • Dialogue History — The chronological record of all spoken (or written) exchanges up to this moment, capturing tone, unresolved tensions, promises, or prior commitments that may influence the character’s strategic choice.
+Subtext and implied meanings are included where relevant.
+  • Interactable Objects — A catalog of physical or digital items within the environment that the character could use, reference, manipulate, or avoid.
+Each object may carry symbolic weight, functional utility, or emotional resonance (e.g., a flickering lantern, a locked diary, a ringing phone).
   • Profile The character’s core personality and background.
@@
 Your evaluation hinges on understanding the nuances of these three cognitive approaches.
-Fast (Intuitive Reaction)
+F.2.1 Fast (Intuitive Reaction)
 A fast response is driven by instinct, emotion, and immediate impressions. It corresponds to Kahneman’s System 1 thinking.
@@
 It should feel like a knee-jerk response, whether clever or foolish.
-Slow (Considered Deliberation)
+F.2.2 Slow (Considered Deliberation)
 A slow response is analytical, rational, and goal-oriented. It corresponds to Kahneman’s System 2 thinking.
@@
 The character is clearly analyzing the situation, managing their impulses, and acting based on a calculated strategy to achieve a specific goal.
-Silence (Strategic Non-Response)
+F.2.3 Silence (Strategic Non-Response)
 Silence is not merely the absence of a response; it is a deliberate, tactical choice.
@@
 The final response is a perfect execution of this deliberate, cool-headed strategy.
-Table 7: An example of labeling the PAD model’s strategic choice. The annotator’s task is to determine the most fitting strategy, which in this case is clearly a Slow Response.
+Table 7: An example of labeling the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: model’s strategic choice.
+The annotator’s task is to determine the most fitting strategy, which in this case is clearly a Slow Response.
 F.4 Important Considerations
@@
   • Quality Assurance: If you are uncertain about the annotation result, please leave it blank rather than making an uncertain guess.
-Appendix G G. Training and Evaluation Details for PAD and HAMLETJudge
+Appendix G Training and Evaluation Details for 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
 G.1 Training Settings and Hyperparameters
-Both PAD and HAMLETJudge models were fine-tuned based on Qwen3-8B using the LLaMA-Factory framework.
+Both 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: models were fine-tuned based on Qwen3-8B using the LLaMA-Factory framework.
 Training was conducted with ZeRO-3 across 8 NVIDIA A800-SXM4-80GB GPUs, utilizing bfloat16 precision.
 The maximum sequence length was set to 65,536, and the total batch size was 128.
-For PAD, we performed supervised fine-tuning (SFT) on the training set for 3 epochs.
+For 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, we performed supervised fine-tuning (SFT) on the training set for 3 epochs.
 The optimizer was configured with a learning rate of 1e-5, a cosine learning rate scheduler, and a warmup ratio of 0.1.
-For HAMLETJudge, we adopted a two-stage training paradigm.
-In the first stage, we conducted supervised fine-tuning on the first half of the training set, using the same optimization setup as PAD (learning rate of 1e-5, cosine scheduler, and 0.1 warmup ratio).
+For 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, we adopted a two-stage training paradigm.
+In the first stage, we conducted supervised fine-tuning on the first half of the training set, using the same optimization setup as 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: (learning rate of 1e-5, cosine scheduler, and 0.1 warmup ratio).
 In the second stage, leveraging the pairwise nature of the dataset, we restructured the remaining half of the training set into chosen–rejected format and applied Direct Preference Optimization (DPO).
@@
 G.2 Data Statistics and Holdout Method
-The complete PAD dataset comprises two main categories: Single Tool Use, which includes Fast, Slow, and Silence responses, and Multi Tool Use, which involves combinations such as Action + Fast, Action + Slow, and Action + Silence.
+The complete 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: dataset comprises two main categories: Single Tool Use, which includes Fast, Slow, and Silence responses, and Multi Tool Use, which involves combinations such as Action + Fast, Action + Slow, and Action + Silence.
 Detailed statistics for each category are presented in Table 8.
-PAD Single Tool Use Multi Tool Use
-Fast Slow Silence Action+Fast Action+Slow Action+Silence
+0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Single Tool Use Multi Tool Use
+Fast Slow Silence
+Action+Fast
+Action+Slow
+Action+Silence
 Data Count 4200 3304 3908 1000 1000 1000
-Mapping Rule speech [thinking]+speech no speech (action)+speech (action)+[thinking]+speech (action)
-Table 8: Statistics of the PAD Dataset.
+Mapping Rule speech
+[thinking]+speech
+no speech
+(action)+speech
+(action)+
+[thinking]+speech
+Table 8: Statistics of the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Dataset.
 The Mapping Rule row in Table 8 illustrates the expected response formats for AI actors under each tool-use strategy.
-The core objective of PAD is to train agents in selecting appropriate response strategies, namely Fast, Slow, or Silence.
+The core objective of 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: is to train agents in selecting appropriate response strategies, namely Fast, Slow, or Silence.
 In multi-tool use settings, the Action Tool is further annotated with a verb-object pair to reflect its parametrized usage in combination with a primary response strategy.
@@
 2) The choice of parameters (i.e., verb and object) is flexible and often has multiple valid options.
-Given these characteristics, we consider the Action Tool as an auxiliary enhancement to the core response strategy in PAD, rather than a primary focus of quantitative evaluation.
+Given these characteristics, we consider the Action Tool as an auxiliary enhancement to the core response strategy in 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, rather than a primary focus of quantitative evaluation.
 The dataset was partitioned via the holdout method (95% training, 5% testing), with stratification to preserve class distribution.
-On the other hand, the HAMLETJudge dataset is designed to evaluate drama performance across three key dimensions: CP (Character Performance), NQ (Narrative Quality), and IE (Interaction Experience).
+On the other hand, the 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: dataset is designed to evaluate drama performance across three key dimensions: CP (Character Performance), NQ (Narrative Quality), and IE (Interaction Experience).
 For each LLM pair—e.g., Claude-3.5-sonnet vs.
 GPT-4.5-preview—75 pairwise comparisons are conducted per dimension in both Chinese and English, resulting in 150 annotated instances per dimension for each model pair.
-We select 7 representative LLMs exhibiting diverse levels of drama performance, as demonstrated on the HAMLET Leaderboard: Claude-4-sonnet-Thinking, DeepSeek-R1-0528, Qwen3-235B-A22B, Gemini-2.5-pro, GPT-4o, Qwen3-8B, and LLaMA-3.1-8B.
+We select 7 representative LLMs exhibiting diverse levels of drama performance, as demonstrated on the 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Leaderboard: Claude-4-sonnet-Thinking, DeepSeek-R1-0528, Qwen3-235B-A22B, Gemini-2.5-pro, GPT-4o, Qwen3-8B, and LLaMA-3.1-8B.
 This leads to a total of $\binom{7}{2}=21$ unique model pairs.
@@
 Detailed statistics are presented in Table 9.
-HAMLETJudge #CP #NQ #IE #Total
+0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: #CP #NQ #IE Total
 Instance Number 3150 3150 3150 9450
-Table 9: Statistics of the HAMLETJudge Dataset.
+Table 9: Statistics of the 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Dataset.
 To ensure fair evaluation, the dataset is splitted using a holdout strategy, with around 95.3% allocated for training and 4.7% for testing.
 Stratified sampling is applied to maintain a balanced distribution across all three dimensions.
-G.3 PAD Latency Penalty
-As discussed in the PAD reliability experiments, although models that output reasoning tokens before invoking tool calls generally achieve better performance, this strategy may introduce intolerable extra latency in real-time drama settings.
-To address this trade-off, we introduce a latency penalty to balance PAD evaluation results.
+G.3 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Latency Penalty
+As discussed in the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: reliability experiments, although models that output reasoning tokens before invoking tool calls generally achieve better performance, this strategy may introduce intolerable extra latency in real-time drama settings.
+To address this trade-off, we introduce a latency penalty to balance 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: evaluation results.
 Accurately and fairly measuring latency is inherently challenging due to several factors: 1) Different model sizes demand varying CUDA memory usage; 2) Certain models, such as DeepSeek-V3 and DeepSeek-R1, are recommended to use FP8 for inference, while most others typically rely on BF16; 3) For closed-source models, latency must be measured via API-level request-response timing, which is subject to various uncontrollable factors.
@@
 Gemini-2.5-pro 21.80 0.10
-PAD 0.36 0
-Table 10: Average Latency per Evaluation Instance and Corresponding Penalty. Note: Penalty is mapped based on latency: 0 for $<$2s, 0.05 for 2–10s, 0.10 for 10–30s, and 0.15 for $\geq$30s.
+0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: 0.36 0
+Table 10: Average Latency per Evaluation Instance and Corresponding Penalty. Note: Penalty is mapped based on latency: 0 for <<2s, 0.05 for 2–10s, 0.10 for 10–30s, and 0.15 for $\geq$30s.
 The average latency results are shown in Table 10.
-Appendix H H. Prompts Used in HAMLET
-This section shows the prompt design in HAMLET multi-agent workflow.
+Appendix H Prompts Used in 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
+This section shows the prompt design used in 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: multi-agent workflow.
 The content between { } is variants or reference templates.
 Responsibilities of all agents and their corresponding prompt designs are as follows:
-H.1. Agents in Offline Planning
+H.1 Agents in Offline Planning
 The Director (Table 11) integrates the narrative structure; the Actor Designer creates the characters (Table 12 and Table 13); the Plot Designer crafts the storyline, environment and interactable props(Table 14 and Table 15).
-H.2. Agents in Online Performance
+H.2 Agents in Online Performance
 During the online performance, agents work together to create a dynamic and coherent drama.
@@
 Finally, the Actor agents (Table 21 and Table 22) bring the characters to life.
-They first use the PAD module to assess the situation and select a response strategy.
+They first use the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: module to assess the situation and select a response strategy.
 Then, they generate the character’s performance, which can be a mix of dialogue, actions, and internal thoughts.
+Appendix I Large Language Models Usage Statement
+During the preparation of this work, we utilized Large Language Models (LLMs), such as ChatGPT, as a writing-assistant tool.
+The use of LLMs was exclusively limited to improving the grammar, clarity, and readability of the text.
+All the scientific ideas, experimental designs, results, and conclusions presented in this paper were conceived and formulated entirely by the human authors.
+The authors take full responsibility for the content of this paper.
 Prompt: Director You are the Director.
@@
 Table 20: The prompt for the Advancer agent to progress the plot.
-Prompt: Actor (PAD Strategy) You are a role-playing expert that can perceive the surrounding environment and decide the appropriate responding strategy to the current speaker under considerable and comprehensive consideration.
+Prompt: PAD module) You are a role-playing expert that can perceive the surrounding environment and decide the appropriate responding strategy to the current speaker under considerable and comprehensive consideration.
 You may call one or more functions to assist with the user query.
@@
 Please decide how to respond with most appropriate tool use or tool use combination.
-Table 21: The prompt for the Actor’s PAD (Perceive, Analyze, Decide) module.
+Table 21: The prompt for the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: module.
 Prompt: Actor (Response Generation) You are role-playing a actor based on the following profile.
@@
 Table 22: The prompt for the Actor agent to generate its final response.
-Appendix I I. Author Contributions
-Sizhou Chen and Shufan Jiang contributed equally to this work.
-Sizhou Chen and Shufan Jiang jointly conceived the study, designed the method, and wrote the manuscript.
-Sizhou Chen prepared the figures.
-Shufan Jiang performed the large-scale model experiments.
-All other authors contributed to the manuscript review.
-All authors have read and agreed to the published version of the manuscript.